### v1

Abstract
We present several equivalent conditions of the continuity of the supremum function $\Sigma C(X)\times\Sigma C(X)\rightarrow\Sigma C(X)$ under mild assumptions, where $C(X)$ denotes the lattice of closed subsets of a $T_{0}$ topological space. We also provide an example of a non-monotone determined space $X$ such that $\eta=\lambda x.{\downarrow}x\colon X\rightarrow\Sigma C(X)$ is continuous. Additionally, we show that a $T_{0}$ space is quasicontinuous (quasialgebraic) iff the lattice of its closed subsets is a quasicontinuous (quasialgebraic) domain by using $n$-approximation. Furthermore, we provide a necessary condition for when a topological space possesses a Scott completion. This allows us to give more examples which do not have Scott completions.
Keywords: Monotone determined space, Quasicontinuous, D-completion, Scott topology, Lattice of closed subsets, Scott completion.
Mathematics Subject Classification: 54A10, 54A20, 06B35.
1 Introduction
Topological semilattice has a deep connection with the theory of continuous lattices in domain theory. For instance, Lawson showed that compact unital semilattices with small semilattices are exactly continuous lattices with Lawson topology, see [11, Theorem VI-3.4]. From the viewpoint of category topology, Banaschewski introduced the notion of an essential embedding in the category of $T_{0}$ spaces and continuous maps [2], which has a close relation to Scott’s injective spaces. He also obtained a characterization of so-called essentially complete $T_{0}$ spaces in terms of join filters. Later Hoffmann provided a more transparent characterization: essentially complete $T_{0}$ spaces are sober spaces which are complete lattices in their specialisation order and satisfy an additional condition connecting the sup-operator with the topology [14]. In this paper we study a special kind of essentially complete $T_{0}$ space, namely the Scott space of the lattice of closed subsets of a topological space with the supremum operator.
Directed spaces [24] are special $T_{0}$ topological spaces, extending the concept of Scott spaces. The idea is that the priori is a $T_{0}$ space and its convergent directed subsets relative to the specialization order rather than a poset and its existing directed supremums. In $T_{0}$ topological spaces, the notion of directed spaces is equivalent to that of monotone determined spaces defined by Erné [8], also is equivalent to crp-spaces introduced by Battenfeld [3, Definition 6.7], in this paper we always assume that topological space is $T_{0}$. The category of directed spaces was shown to be Cartesian closed [24, 20]. $T_{0}$ c-spaces (or $n$-continuous spaces) are all monotone determined spaces. Based on this, Feng and Kou [9] introduced another approximation relation on monotone determined spaces, called $d$-approximation, and defined the quasicontinuous monotone determined spaces. They established an equivalence between quasicontinuous spaces and locally hypercompact spaces. Recently, Zhang, Shi and Li [25] have also contributed to this field by introducing the notions of continuous spaces and quasicontinuous spaces, focusing on these spaces within weak monotone convergence spaces.
It is well known that the hull-kernel topology of the spectrum ${\mathrm{Spec}L}$ equals to the Scott topology when $L$ is a completely distributive (or distributive hypercontinuous) lattice, see [11]. There is a long-standing open problem asked by Lawson and Mislove that for which distributive continuous lattice its spectrum with the hull-kernel topology is exactly a sober locally compact Scott space (see [23, Problem 528]). Chen, Kou and Lyu [4] found a necessary condition for this problem, which is $\sigma(L^{op})=v(L^{op})$. Since $L^{op}$ can be viewed as a lattice of closed subsets of a topological space, we continue to investigate the equivalent conditions of $\sigma(C(X))=v(C(X))$ for general topological spaces $X$. Surprisingly the continuity of the supremum function $\Sigma C(X)\times\Sigma C(X)\rightarrow\Sigma C(X)$ is one of these equivalent conditions under mild assumptions. For monotone determined spaces, we can drop the extra assumption. In addition, we provide an example of non-monotone determined space $X$ such that $\eta=\lambda x.{\downarrow}x\colon X\rightarrow\Sigma C(X)$ is continuous.
In domain theory, there is a close correspondence between the continuity of a dcpo $L$ and the continuity of the lattice of its Scott open subsets $\sigma(L)$. More surprisingly, the continuity of a dcpo $L$ can be characterized by the continuity of its lattice of its Scott closed subsets $\Gamma(L)$ (for example, see [4, Theorem 3.17]). These correspondences can be extended to general $T_{0}$ spaces. We show that a $T_{0}$ space $X$ is quasicontinuous (quasialgebraic) iff the lattice of its closed subsets $C(X)$ is quasicontinuous (quasialgebraic).
Monotone convergence spaces are very important in domain theory. Every $T_{0}$ space has a free monotone convergence completion called D-completion (for example, see [16]). We show that D-completion of any monotone determined space is a Scott space. The concept of Scott completion is introduced by Zhang, Shi and Li in [25], which has a slight difference with D-completion. We also obtain that Scott-completion of any monotone determined space coincides with its D-completion as a corollary. The Scott-completion of monotone determined spaces are shown to be very useful to construct free dcpo algebras over any general dcpo (see [6, 7]). But the notion does not seem to have been studied much. We provide a necessary condition for when a space has Scott completion. Based on this we give a simple proof of a previous result in [17, Theorem 4.1].
Outline. In Section 2, we recall some useful knowledge about monotone determined spaces. We devote Section 3 to a discussion of D-completion of monotone determined spaces. In Section 4 the Scott space of the lattice of closed subsets of a topological space with the jointly continuous supremum operator is discussed. We characterize the quasicontinuity of lattice of closed subsets in Section 5. Finally more examples of topological spaces which do not have Scott completions are presented in Section 6.
2 Preliminaries
We assume some basic knowledge of domain theory and topology, as in, e.g., [1, 11, 12].
A nonempty set $P$ endowed with a partial order $\leq$ is called a poset. For $A\subseteq P$, we set $\downarrow A=\{x\in P:\exists a\in A,\ x\leq a\}$$\uparrow A=\{x\in P:\exists a\in P,\ a\leq x\}$$A$ is called a lower or an upper set, if $A=\downarrow A$ or $A=\uparrow A$ respectively. For an element $a\in P$, we use $\downarrow a$ or $\uparrow a$ instead of $\downarrow\{a\}$ or $\uparrow\{a\}$, respectively.
Let $P$ be a poset. We denote $\sigma(P)$, $\upsilon(P)$ and $A(P)$ to be the Scott topology, the upper topology and the Alexandroff topology on $P$ respectively. For a complete lattice $L$, and $x,y\in L$, we say that $x\prec y$ iff $y\in\mathrm{int}_{\upsilon(L)}\uparrow x$. $L$ is hypercontinuous if $\{d\in L:d\prec x\}$ is directed and has $x$ as a supremum for all $x\in L$. $L$ is hyperalgebraic if $x=\bigvee\{d\in L,d\prec d\leq x\}$ for all $x\in L$. An element $y$ such that $y\prec y$ is called hypercompact.
[…]
For any two topological spaces $X,Y$, we denote $Y^{X}$ or $TOP(X,Y)$ the set of all continuous maps from $X$ to $Y$, endowed with the pointwise order. Denote $[X\to Y]_{p}$ and $[X\to Y]_{I}$ to be the topological space equipped with the topology of pointwise convergence and the Isbell topology on $Y^{X}$ respectively.
We now introduce the notion of a monotone determined space.
Let $(X,\mathcal{O}(X))$ be a $T_{0}$ space. Every directed subset $D\subseteq X$ can be regarded as a monotone net $(d)_{d\in D}$. Set $DS(X)=\{D\subseteq X:D\ {\rm is\ directed}\}$ to be the family of all directed subsets of $X$. For an $x\in X$, we denote $D\rightarrow x$ to mean that $x$ is a limit of $D$, i.e., $D$ converges to $x$ with respect to the topology on $X$. Then the following result is obvious.
Lemma 2.1.
Let $X$ be a $T_{0}$ space. For any $(D,x)\in DS(X)\times X$, $D\rightarrow x$ if and only if $D\cap U\not=\emptyset$ for any open neighborhood of $x$.
Set $DLim(X)=\{(D,x)\in DS(X)\times X:\ D\rightarrow x\}$ to be the set of all pairs of directed subsets and their limits in $X$. Then ($(\{y\},x)\in DLim(X)$ iff $x\sqsubseteq y$ for all $x,y\in X$.
Definition 2.2.
Let $X$ be a $T_{0}$ space. A subset $U\subseteq X$ is called directed-open if for all $(D,x)\in DLim(X)$, $x\in U$ implies $D\cap U\not=\emptyset$.
Obviously, every open set of $X$ is directed-open. Set $d(\mathcal{O}(X))=\{U\subseteq X:U\ {\rm is}\ {\rm directed}\text{-}{\rm open}\},$ then $\mathcal{O}(X)\subseteq d(\mathcal{O}(X))$.
Theorem 2.3.
[24] Let $X$ be a $T_{0}$ topological space. Then
  1. (1)
    For all $U\in d(\mathcal{O}(X))$, $U=\uparrow U$;
  2. (2)
    $X$ equipped with $d(\mathcal{O}(X))$ is a $T_{0}$ topological space such that $\sqsubseteq_{d}=\sqsubseteq$, where $\sqsubseteq_{d}$ is the specialization order relative to $d(\mathcal{O}(X))$.
  3. (3)
    For a directed subset $D$ of $X$, $D\rightarrow x$ iff $D\rightarrow_{d}x$ for all $x\in X$, where $D\rightarrow_{d}x$ means that $D$ converges to $x$ with respect to the topology $d(\mathcal{O}(X))$.
  4. (4)
    $d(d(\mathcal{O}(X))=d(\mathcal{O}(X))$.
Definition 2.4.
[24] A topological space $X$ is said to be a monotone determined space if it is $T_{0}$ and every directed-open set is open; equivalently, $d(\mathcal{O}(X))=\mathcal{O}(X)$.
Given any space $X$, we denote $\mathcal{D}X$ the topological space $(X,d(\mathcal{O}(X)))$.
Theorem 2.5.
[24] Let $X$ be a $T_{0}$ space, we have
  1. (1)
    $\mathcal{D}X$ is a monotone determined space.
  2. (2)
    The following three conditions are equivalent to each other:
    1. (i)
      $X$ is a monotone determined space;
    2. (ii)
      For all $U\subseteq X$, $U$ is open iff for any $(D,x)\in DLim(X)$, $x\in U$ implies $U\cap D\not=\emptyset$.
    3. (iii)
      For all $A\subseteq X$, $A$ is closed iff for any directed subset $D\subseteq A$, $D\rightarrow x$ implies $x\in A$ for all $x\in X$.
Monotone determined spaces include many important examples in domain theory.
Example 2.6.
  1. (1)
    Every poset endowed with the Scott topology is a monotone determined space.
  2. (2)
    ([8]) Every poset endowed with the weak Scott topology is a monotone determined space.
  3. (3)
    Every c-space is a monotone determined space. In particular, any poset endowed with the Alexandroff topology is a monotone determined space.
  4. (4)
    ([8, 9]) Every locally hypercompact space is a monotone determined space.
We denote ${\bf DTop}$ the category of all nonempty monotone determined spaces with continuous maps as morphisms.
Lemma 2.7.
Given any two monotone determined spaces $X,Y$, a subset $U$ of $X\otimes Y$ is open iff the following conditions hold:
  1. (1)
    for any directed subsets $(x_{i})_{I}$ with $(x_{i})_{I}\to x$ in $X$ and any $(x,y)\in U$, we have $((x_{i},y))_{I}\cap U\neq\emptyset$;
  2. (2)
    for any directed subsets $(y_{i})_{I}$ with $(y_{i})_{I}\to y$ in $Y$ and any $(x,y)\in U$ we have $((x,y_{i}))_{I}\cap U\neq\emptyset$.
Proof. Assume that $U$ is open in $X\otimes Y$ and $(x,y)\in U$. For any directed set $(x_{i})_{I}\to x$ in $X$ and any $y\in Y$, $((x_{i},y))_{I}$ is a directed subset of $X\otimes Y$ and $((x_{i},y))_{I}\to(x,y)$ in $X\times Y$. By Theorem 2.3 (3), $((x_{i},y))_{I}\to(x,y)$ in $X\otimes Y$. Thus, $((x_{i},y))_{I}\cap U\neq\emptyset$. It is the same for $(y_{i})_{I}$.
Conversely, assume that $(1)$ and $(2)$ are satisfied. We show that $U$ is open in $X\otimes Y$. It is easily seen that $U$ is an upper set relative to the specialization order of $X\times Y$. Let $D=((x_{i},y_{i}))_{I}$ be a directed subset of $X\times Y$ and converges to $(x,y)\in U$ in $X\times Y$. We have that $(x_{i})_{I}\to x$ in $X$ and $(y_{i})_{I}\to y$ in $Y$ respectively. Thus, $((x_{i},y))_{I}\to(x,y)$ in $X\times Y$, and then there exists some $i_{0}\in I$ such that $(x_{i_{0}},y)\in U$. By $((x_{i_{0}},y_{i}))_{I}\to(x_{i_{0}},y)$, there exists some $i_{1}\in I$ such that $(x_{i_{0}},y_{i_{1}})\in U$. Let $i_{0},i_{1}\leq i_{2}$, then $(x_{i_{2}},y_{i_{2}})\in U$. $\Box$
3 D-completion of monotone determined spaces
Monotone convergence spaces (a.k.a d-spaces) is a very important class of topological spaces in domain theory. It is well known that every $T_{0}$ space has a D-completion. Let us recall some results about D-completion for further discussions.
Definition 3.8.
Let $X$ be a $T_{0}$ space. If $X$ with the specialization order is a dcpo and every open set of $X$ is Scott open in $(X,\leq)$, then we call $X$ a monotone convergence space space.
A subset $A$ of a poset $P$ is called d-closed if for every directed set $D$ of $A$ that possesses supremum $\bigvee D$, $\bigvee D$ is included in $A$. The d-closed sets form the closed sets for a topology, called the d-topology [16]. We denote the closure of $A$ in $P$ with d-topology by $\text{cl}_{d}(A)$. If $\text{cl}_{d}(A)$ is equal to $P$, then we say that $A$ is d-dense in $P$. A map $f$ from a poset $P$ to a poset $Q$ is called d-continuous if $f$ is continuous under the d-topology.
Definition 3.9.
[16] An embedding $j\colon X\rightarrow\tilde{X}$ of a space $X$ with image a d-dense subset of a d-space $\tilde{X}$ is called a $D$-completion.
Lemma 3.10.
[16] Consider the following properties for a subset $A$ of a $T_{0}$ space
  1. (1)
    $A$ is a monotone convergence space.
  2. (2)
    $A$ is a sub-dcpo.
  3. (3)
    $A$ is d-closed.
Then $(1)$ implies $(2)$ implies $(3)$, and all three are equivalent if $X$ is a monotone convergence space.
The following result says that the D-completions are universal and hence they are unique up to isomorphism.
Theorem 3.11.
[16] Let $j\colon X\rightarrow Y$ be a topological embedding of a space $X$ into a monotone convergence space $Y$. Let $\tilde{X}=\text{cl}_{d}(j(X))$ be the d-closure of $j(X)$ in $Y$, equipped with the relative topology from $Y$. Then $k\colon X\rightarrow\tilde{X}$, the corestriction of $j$, is a universal $D$-completion, that is, for every continuous map $f$ from $X$ to a monotone convergence space $M$, there is a unique continuous map $\tilde{f}\colon\tilde{X}\rightarrow M$ such that $\tilde{f}\circ k=f$.
Remark 3.12.
For any $T_{0}$ space $X$, there is a natural monotone convergence space $C(X)$ (the lattice of closed subsets of $X$) with the upper topology such that $\eta=\lambda x.{\downarrow}x\colon X\rightarrow C(X)$ is an embedding. We call the d-closure of $\eta(X)$ in $C(X)$ the standard D-completion of $X$ and denote it by $X^{d}$.
Proposition 3.13.
Let $X$ be a $T_{0}$ space and $X^{d}$ be its D-completion. Then $\mathcal{O}(X)\cong\mathcal{O}(X^{d})$.
Proof.
Let $f\colon X\rightarrow X^{s}$ be the soberification map. Since $X^{s}$ is a d-space, the D-completion $X^{d}$ of $X$ can be regarded as a subspace of $X^{s}$ by Theorem 3.11, then $X^{s}$ is also a soberification of $X^{d}$. We conclude that $\mathcal{O}(X)\cong\mathcal{O}(X^{d})$ directly. ∎
Theorem 3.14.
Let $X$ be a monotone determined space and $X^{d}$ be the D-completion of $X$. Then $X^{d}$ is a Scott space, i.e. the topology on $X^{d}$ equals to the Scott topology on $(X^{d},\leq)$.
Proof.
Since $X^{d}$ is a monotone convergence space, we only need to show that every Scott open set of $(X^{d},\leq)$ is open in $X^{d}$. For any Scott open set $\mathcal{U}$ of $(X^{d},\leq)$ and any closed set $A\in\mathcal{U}$, notice that
[…]
Remark 3.15.
The aforementioned theorem also appeared in [3, Proposition 6.13] with a different proof.
The following results are direct consequences of the above theorem.
Corollary 3.16.
  1. (1)
    [16, Theorem 7.4] Let $P$ be a poset. The D-completion of $\Sigma P$ is a Scott space.
  2. (2)
    [25, Theorem 4.7] Let $X$ be a monotone determined space. $(X^{d},i)$ is a Scott completion of $X$.
The D-completion of any monotone determined space is a Scott space while there exists a non-monotone determined space whose D-completion is a Scott space. This example also can be found in [3].
Example 3.17.
Let $D$ be a set ×$\mathbb{N}\times\mathbb{N}\cup\{a_{n}:n\in\mathbb{N}\}\cup\{\top\}$. The order on $D$ is defined as follows: $x\leq y$ iff
  • 1.
    $y=\top$;
  • 2.
    There are natural numbers $n\leq m$ such that $x=a_{n},y=a_{m}$;
  • 3.
    There are natural numbers $n$ and $m_{1}\leq m_{2}$ such that $x=(n,m_{1}),y=(n,m_{2})$;
  • 4.
    There are natural numbers $n$ and $m$ such that $x=(n,m),y=a_{n}$.
[…]
$a_{1}$$a_{2}$$a_{3}$$a_{3}$$\top$
Figure 1: ×$\mathbb{N}\times\mathbb{N}\cup\{a_{n}:n\in\mathbb{N}\}\cup\{\top\}$
4 Continuity of the supremum function $\Sigma C(X)\times\Sigma C(X)\rightarrow\Sigma C(X)$
It is well known that the spectrum with hull-kernel topology of a completely distributive lattice (resp. a distributive hypercontinuous lattice) is exactly a continuous (resp. quasicontinuous) dcpo endowed with the Scott topology [10, 15, 18]. There is a long-standing open problem asked by Lawson and Mislove that, for which distributive continuous lattice its spectrum is exactly a sober locally compact Scott space (see [23, Problem 528]). Chen, Kou and Lyu [4] found a necessary condition for this problem, which is $\sigma(L^{op})=v(L^{op})$. Since $L^{op}$ can be viewed as a lattice of closed subsets of a topological space, we continue to investigate the equivalent conditions of $\sigma(C(X))=v(C(X))$ for general topological spaces $X$. Surprisingly the continuity of the supremum function $\Sigma C(X)\times\Sigma C(X)\rightarrow\Sigma C(X)$ is one of these equivalent conditions under mild assumptions.
Given any poset $P$, $\upsilon(\Gamma(P))=\sigma(\Gamma(P))$ is a necessary condition for $\Sigma P$ to be core compact [4]. Given a topological space $X$, we denote $\prod\limits^{n}X$ to be topological product of $n$ copies of $X$. For any $n\in\mathbb{N}$, define a map $s_{n}:\prod\limits^{n}X\to\Sigma(C(X))$ as follows: $\forall(x_{1},x_{2},\dots,x_{n})\in\prod\limits^{n}X$,
$s_{n}(x_{1},x_{2},\dots,x_{n})=\ \downarrow\!{\{x_{1},x_{2},\dots,x_{n}\}}.$
The following result is due to R.Hoffmann.
Theorem 4.18.
[14, Proposition 1.1] A space is an essentially complete $T_{0}$ space iff the following are satisfied:
  1. (i)
    The associated pre-order $(X,\leq)$ is a complete lattice.
  2. (ii)
    The topology on $X$ is coarser than the Scott topology of $(X,\leq)$.
  3. (iii)
    Whenever $a=\sup\{b,c\}$ with $a,b,c\in X$, then $\{U\cap V|b\in U\in\mathcal{O}(X),c\in U\in\mathcal{O}(X)\}$ is an open neighborhood basis for $a$ in $X$.
Hoffmann noticed that the above condition $(iii)$ can be replaced by that “$\sup\colon X\times X\rightarrow X$ is continuous” [14, Lemma 1.5].
Proposition 4.19.
Let $X,Y,Z$ be monotone determined spaces. A map $f:X\otimes Y\to Z$ is continuous iff it is separately continuous.
Theorem 4.20.
For a topological space $X$, the following conditions are equivalent:
  • (1)
    $\sigma(C(X))=\upsilon(C(X))$;
  • (2)
    $s_{n}$ is continuous for all $n\in\mathbb{N}$;
  • (3)
    $s_{1}$ is continuous and $\bigcup\colon\Sigma C(X)\times\Sigma C(X)\rightarrow\Sigma C(X)$ is continuous.
Proof.
$(1)\implies(3)$. Notice that
$\Diamond U=\{A\in C(X):A\cap U\neq\emptyset\},U\in\mathcal{O}(X)$
form a subbasis of the upper topology on $C(X)$. It is a direct argument.
$(3)\implies(2)$. From
$X\times X\stackrel{{\scriptstyle s_{1}\times s_{1}}}{{\longrightarrow}}\Sigma C(X)\times\Sigma C(X)\stackrel{{\scriptstyle\bigcup}}{{\longrightarrow}}\Sigma C(X),$
we obtain that $s_{2}=\bigcup\circ(s_{1}\times s_{1})$. It follows that $s_{2}$ is continuous. By induction on $n$, it is easy to see that $s_{n}$ is continuous.
$(2)\implies(1)$. For the completeness, we give the detailed proof here. Assume that $s_{n}$ is continuous for all $n\in\mathbb{N}$. Let $\mathcal{U}$ be an open subset of $\Sigma(C(X))$ and $A\in\mathcal{U}$. Without loss of generality, we assume $A\not=\emptyset$. Note that since $A=\bigcup\{\downarrow\!F:F\subseteq_{f}\!A\}$ and $\{\downarrow F:F\subseteq_{f}A\}$ is a directed family in $C(X)$, there exists a non-empty finite subset $F$ of $A$ such that $\downarrow\!F\in\mathcal{U}$. Let $F=\{x_{1},x_{2},\ldots,x_{n}\}$, then $s_{n}(x_{1},x_{2},\ldots,x_{n})=\downarrow\!F\in\mathcal{U}$. It follows that $(x_{1},x_{2},\ldots,x_{n})\in s_{n}^{-1}(\mathcal{U})$. By the continuity of $s_{n}$, there exists a family of open subsets $U_{k}(1\leq k\leq n)$ such that $U_{1}\times U_{2}\times\cdots\times U_{n}$ is open in $\prod\limits^{n}X$ and
$(x_{1},x_{2},\ldots,x_{n})\in U_{1}\times U_{2}\times\cdots\times U_{n}\subseteq s_{n}^{-1}(\mathcal{U}).$
Since $x_{k}\in A$ for $1\leq k\leq n$, we have $A\in\Diamond U_{k}=\{B\in C(X):\ B\cap U_{k}\not=\emptyset\}$. It follows that $A\in\bigcap\limits_{k=1}^{n}\Diamond U_{k}\in\upsilon(C(X))$. For any $B\in\bigcap\limits_{k=1}^{n}\Diamond U_{k}$, there exists $y_{k}\in B\cap U_{k}$ for $1\leq k\leq n$. Since $(y_{1},y_{2},\dots,y_{n})\in s_{n}^{-1}(\mathcal{U})$, we have $\bigcup\limits_{k=1}^{n}\downarrow y_{k}\in\mathcal{U}$. It follows that $B\in\mathcal{U}$, i.e., $A\in\bigcap\limits_{k=1}^{n}\Diamond U_{k}\subseteq\mathcal{U}$. ∎
Remark 4.21.
The equivalence of $(1)$ and $(2)$ has already been given in [5, Proposition 4.1].
[…]
Figure 2: P
Example 4.22.
There exists a space $X$ such that $\bigcup\colon\Sigma C(X)\times\Sigma C(X)\rightarrow\Sigma C(X)$ is continuous while $s_{1}\colon X\rightarrow\Sigma C(X)$ is not continuous. We consider the example $P$ with its upper topology from [4, Example 4.9], let $X=(P,v(P))$. It is easy to see that $C(X)=\{{\downarrow}F:F\subseteq_{fin}P\}\cup\{P,\emptyset\}$. Next we verify that $\bigcup\colon\Sigma C(X)\times\Sigma C(X)\rightarrow\Sigma C(X)$ is continuous. For any Scott open subset $\mathcal{U}$ of $C(X)$, $\bigcup^{-1}(\mathcal{U})=\{(A,B)\in C(X)\times C(X):A\cup B\in\mathcal{U}\}$. For any $A\cup B\in\mathcal{U}$, we construct two subsets of $C(X),\mathcal{V}=\mathcal{U}\cup\{A^{\prime}\in C(X):A\subseteq A^{\prime}\}$, $\mathcal{W}=\mathcal{U}\cup\{B^{\prime}\in C(X):B\subseteq B^{\prime}\}$. We claim that both $\mathcal{V}$ and $\mathcal{W}$ are Scott open subsets of $C(X)$. Obviously $\mathcal{V}$ is upward closed. For any directed family $\{F_{i}\}_{i\in I}$ consisting of closed subsets of $X$ such that $\sup^{\uparrow}F_{i}\in\mathcal{V}$, if $\sup^{\uparrow}F_{i}=P\in\mathcal{V}$, then there exists some $F_{i}\in\mathcal{U}\subseteq\mathcal{V}$ since $\mathcal{U}$ is Scott open. If $\sup^{\uparrow}F_{i}\neq P$, then there is a maximal element $F_{i_{0}}$ of $\{F_{i}:i\in I\}$. Hence $\mathcal{V}$ is Scott open. It can be deduced that $\mathcal{W}$ is Scott open as well. Notice that $(A,B)\in\mathcal{V}\times\mathcal{W}\subseteq\bigcup^{-1}(\mathcal{U})$. Therefore $\bigcup\colon\Sigma C(X)\times\Sigma C(X)\rightarrow\Sigma C(X)$ is continuous. Since $\sigma(C(X))\neq v(C(X))$ from [4, Example 4.9], we conclude that $s_{1}\colon X\rightarrow\Sigma C(X)$ is not continuous by Theorem 4.20.
Corollary 4.23.
Let $X$ be a monotone determined space. The following statements are equivalent:
  • (1)
    $\sigma(C(X))=\upsilon(C(X))$;
  • (2)
    $s_{n}$ is continuous for all $n\in\mathbb{N}$;
  • (3)
    $\bigcup\colon\Sigma C(X)\times\Sigma C(X)\rightarrow\Sigma C(X)$ is continuous.
Proof.
Notice that $s_{1}\colon X\rightarrow\Sigma C(X)$ is continuous since $X$ is a monotone determined space. ∎
Remark 4.24.
Let $X$ be a monotone determined space. We consider a topology $\tau$ on $C(X)$ with $v(C(X))\subseteq\tau\subseteq\sigma(C(X))$. By a similar argument in Theorem 4.20, we can show that $\tau$ is equal to $v(C(X))$ if the operator $\bigcup\colon(C(X),\tau)\times(C(X),\tau)\rightarrow(C(X),\tau)$ is continuous. In other words, there exists only one topology $\tau$ on $C(X)$ such that $(1)$ the specialization order of $\tau$ is equal to the inclusion order and $(2)$ $(C(X),\tau)$ is an essentially complete $T_{0}$ space. But this is false for a non-monotone determined space $X$, for instance see Example 4.22.
Corollary 4.25.
Let $D$ be a poset. The following conditions are equivalent:
  • (1)
    $\sigma(\Gamma(D))=\upsilon(\Gamma(D))$;
  • (2)
    $s_{n}$ is continuous for all $n\in\mathbb{N}$;
  • (3)
    $\bigcup\colon\Sigma\Gamma(D)\times\Sigma\Gamma(D)\rightarrow\Sigma\Gamma(D)$ is continuous.
The following result is a direct consequence of the above corollary.
Corollary 4.26.
[4, Proposition 3.5] Let $P$ be a poset. If ${\displaystyle\Sigma(\prod^{n}P)=\prod^{n}(\Sigma P)}$ for each $n\in\mathbb{N}$, then $\sigma(\Gamma(P))=v(\Gamma(P))$.
Proposition 4.27.
Let $X$ be a $T_{1}$ space. The following conditions are equivalent:
  1. (1)
    $X$ has the discrete topology;
  2. (2)
    $\sigma(C(X))=\upsilon(C(X))$;
  3. (3)
    $s_{1}\colon X\rightarrow\Sigma C(X)$ is continuous.
Proof.
$(1)\implies(2)$. Since $C(X)$ is a completely distributive lattice, $\sigma(C(X))=\upsilon(C(X))$.
$(2)\implies(3)$. Straightforward.
$(3)\implies(1)$. Notice that $\mathcal{A}=\{\{a\}:a\in A\}\cup\{\emptyset\}$ is always a Scott closed set in $C(X)$ for any $A\subseteq X$. Take the inverse image of $\mathcal{A}$, we obtain that $s_{1}^{-1}(\mathcal{A})=A$ is closed in $X$ since $\eta$ is continuous. Hence $X$ has the discrete topology. ∎
Remark 4.28.
$(2)$ implies $(1)$ can be found in [4, Proposition 3.15].The above results does not hold for $T_{0}$ spaces. For example we consider $X=\Sigma L$ where $L$ is the Isbell’s lattice, then $(3)$ holds while $(2)$ does not hold since $\Sigma C(X)$ is not sober [21].
A natural question arises that whether a topological space $X$ that makes the map $s_{1}\colon X\rightarrow\Sigma(C(X))$ continuous is a monotone determined space? When $L$ is a monotone convergence space such that $(L,\leq)$ is a complete lattice, the answer is positive (see [5, Proposition 4.11]). But this is not true for general topological spaces. First we recall a useful result about the product topology of Scott spaces.
Proposition 4.29.
[22, Lemma 4.1] Let $P,Q$ be two posets. If $|\text{Id}(P)|$, $|\text{Id}(Q)|$ are both countable, then $\Sigma(P\times Q)=\Sigma P\times\Sigma Q$. ($\text{Id}(P)$ means the set of all non-principal ideals of $P$.)
Example 4.30.
There exists a non-monotone determined space $X$ such that $\eta\colon X\rightarrow\Sigma C(X)$ is continuous and $\sigma(C(X))=v(C(X))$. We consider the space $X$ in Example 3.17, it follows that $C(X)\cong\Gamma(D)$ since $\Sigma D$ is the D-completion of $X$. Notice that ${\displaystyle Id(\prod^{n}D)=\prod^{n}Id(D)}$, by using Corollary 4.26 and Proposition 4.29 we obtain that $\sigma(\Gamma(D))=v(\Gamma(D))$. Hence we also have $\sigma(C(X))=v(C(X))$. It is easy to see that $\eta\colon X\rightarrow\Sigma(C(X))$ is continuous.
Next, we consider the lattice of Scott closed subsets of complete lattices. The following lemma is an equivalent description of [4, Proposition 4.10].
Lemma 4.31.
Let $L,M$ be complete lattices and $S=L\times M$. If ${}\bigvee\colon\Sigma S\times\Sigma S\rightarrow\Sigma S$ is continuous, then $\Sigma(L\times M)=\Sigma L\times\Sigma M$.
By the above lemma, if $\Sigma L\times\Sigma M$ is not a Scott space, then ${}\bigvee\colon\Sigma S\times\Sigma S\rightarrow\Sigma S$ is not continuous. We recall that a topological space $X$ is called a Scott space if its topology is equal to the Scott topology on $(X,\leq)$, which $\leq$ is the specialization order of $X$.
Proposition 4.32.
[11, Proposition II-3.15(iii)] If $X$ is a continuous retract of a monotone convergence space $Y$, then $X$ must be $T_{0}$ and is a monotone convergence space. In the special case that $Y$ is a dcpo equipped with the Scott topology, then $X$ must also have the Scott topology.
Theorem 4.33.
Let $L$ be a complete lattice. Consider the following conditions
  1. (1)
    $\bigcup\colon\Sigma\Gamma(L)\times\Sigma\Gamma(L)\rightarrow\Sigma\Gamma(L)$ is continuous;
  2. (2)
    $\bigvee\colon\Sigma L\times\Sigma L\rightarrow\Sigma L$ is continuous.
Then $(1)\implies(2)$, but the converse direction does not hold generally.
Proof.
$(1)\implies(2)$.
Notice that both $\eta=\lambda x.{\downarrow}x\colon L\rightarrow\Gamma(L)$ and $\theta=\lambda A.\sup A\colon\Gamma(L)\rightarrow L$ are Scott continuous. Clearly $\theta\circ\eta=1_{L}$. Hence $\Sigma L$ is a continuous retraction of $\Sigma\Gamma(L)$. It is direct to verify that $\bigvee=\sup\circ\bigcup\circ(\eta\times\eta)\colon\Sigma L\times\Sigma L\rightarrow\Sigma L$. So $\bigvee\colon\Sigma L\times\Sigma L\rightarrow\Sigma L$ is continuous.
We use Hertling’s example [13, Theorem 7.4] to show that there exists a complete lattice $Z$ such that condition $(2)$ holds but condition $(1)$ fails. $Z$ is the disjoint union of a complete lattice $X$ and $Y=\sigma(X)$ plus a top element and a bottom element, i.e., $Z=\{X+Y\}\bigcup\{\top,\bot\}$. $Z$ has the following conditions:
  • 1.
    $\Sigma X\times\Sigma Y\neq\Sigma(X\times Y)$, $\Sigma Z\times\Sigma Z\neq\Sigma(Z\times Z)$.
  • 2.
    $\bigvee\colon\Sigma Z\times\Sigma Z\rightarrow\Sigma Z$ is continuous.
Assume that $\bigcup\colon\Sigma\Gamma(Z)\times\Sigma\Gamma(Z)\rightarrow\Sigma\Gamma(Z)$ is continuous. It is easy to see that $\Gamma(Z)\cong\Gamma(X)\times\Gamma(Y)\bigcup\{\top,\bot\}$, the isomorphic map $\theta\colon\Gamma(Z)\rightarrow\Gamma(X)\times\Gamma(Y)\bigcup\{\top,\bot\}$ is defined as follows: $\theta(Z)=\top,\theta(\emptyset)=\bot,\theta(A)=(A\cap X,A\cap Y)$ if $A$ is a proper nonempty Scott closed subset of $Z$.
Clearly $\Sigma(\Gamma(X)\times\Gamma(Y))$ is a subspace of $\Sigma\Gamma(Z)$. Hence $\bigvee\colon\Sigma(\Gamma(X)\times\Gamma(Y))\times\Sigma(\Gamma(X)\times\Gamma(Y))\rightarrow\Sigma(\Gamma(X)\times\Gamma(Y))$ is continuous. By Lemma 4.31, $\Sigma(\Gamma(X)\times\Gamma(Y))=\Sigma\Gamma(X)\times\Sigma\Gamma(Y)$. Notice that $\Sigma X\times\Sigma Y$ is a retraction of $\Sigma\Gamma(X)\times\Sigma\Gamma(Y)$ which is a Scott space. It follows that $\Sigma X\times\Sigma Y$ is also a Scott space by Proposition 4.32, i.e., $\Sigma X\times\Sigma Y=\Sigma(X\times Y)$. A contradiction. Consequently we also have $\sigma(\Gamma(Z))\neq v(\Gamma(Z))$.
5 Quasicontinuity of the lattice of closed subsets
We introduce the concepts of quasicontinuous spaces and quasialgebraic spaces. Then we will summarize the correspondences among the continuity of a $T_{0}$ space, the continuity of the lattice of its open subsets, and the continuity of the lattice of its closed subsets.
Given any set $X$, we use $F\subseteq_{f}X$ to denote that $F$ is a finite subset of $X$. Given any two subsets $G,H\subseteq X$, we define $G\leq H$ iff $\uparrow H\subseteq\uparrow G$. A family of finite sets is said to be directed if given $F_{1},F_{2}$ in the family, there exists an $F$ in the family such that $F_{1},F_{2}\leq F$. Let $X$ be a $T_{0}$ topological space and $\mathcal{M}$ be a directed family of finite subsets of $X$. Then $F_{\mathcal{M}}=\{A\subseteq X:\exists M\in\mathcal{M},\uparrow M\subseteq A\}$ is a filter under the inclusion order. We say that $\mathcal{M}$ converges to $x$ if $F_{\mathcal{M}}$ converges to $x$, i.e., for any open neighbourhood $U$ of $x$, there exists some $A\in\mathcal{M}$ such that $A\subseteq U$.
Definition 5.34.
Let $X$ be a $T_{0}$ space and $G,H\subseteq X$. We say that $G$ $n$-$approximates$ $H$, denoted by $G\ll_{n}H$, if for every net $(x_{j})_{J}$ of $X$, $(x_{j})_{J}\to y$ for some $y\in H$ implies $(x_{j})_{J}\cap\uparrow G\neq\emptyset$. We write $G\ll_{n}x$ for $G\ll_{n}\{x\}$. $G$ is said to be $n$-compact if $G$ is a finite set and $G\ll_{n}G$.
Lemma 5.35.
For any $T_{0}$ space $X$, the following are equivalent: for all $x,y\in Y$,
  1. (1)
    $G\ll_{n}H$;
  2. (2)
    $H\subseteq(\uparrow G)^{\circ}$, where $(\uparrow G)^{\circ}$ means the interior of $\uparrow G$.
Proof. $(2)\Rightarrow(1)$ is obvious. We only show $(1)\Rightarrow(2)$. Assume that $y\in H,y\not\in(\uparrow G)^{\circ}$. Then, for any open set $U\subseteq X$ with $y\in U$, there exists a $y_{U}\in U$ such that $y_{U}\not\in\uparrow G$. Set $J=\{U\in\mathcal{O}(X):y\in U\}$ ordered as: $\forall U,V\in J$,$U\leq V\Leftrightarrow V\subseteq U.$ Then $J$ is directed. Easily, one can see that $(y_{U})_{U\in J}$ is a net converging to $y$. Since $G\ll_{n}y$, there exists $U_{0}\in J$ such that $x\sqsubseteq y_{U_{0}}$ for some $x\in G$ by definition. It is a contradiction for $y_{U_{0}}\not\in\uparrow G$. Hence, $y\in(\uparrow G)^{\circ}$.  $\Box$
Since the $n$-approximation does not depend on monotone determined space, a notion of $n$-quasicontinuity for a $T_{0}$ space can be defined as follows.
Definition 5.36.
A $T_{0}$ topological space $X$ is called $n$-quasicontinuous if for each $x\in X$, there is a directed family $D(x)\subseteq fin_{n}(x)=\{F:F\ \text{is finite},F\ll_{n}x\}$ such that $D(x)\rightarrow x$.
Definition 5.37.
A $T_{0}$ topological space $X$ is called locally hypercompact (or finitary compact), if for any open subsets $U$ and $x\in U$, there exists some $F\subseteq_{f}X$ such that $x\in(\uparrow F)^{\circ}\subseteq\uparrow F\subseteq U$. $X$ is called hypercompactly based if for any open subsets $U$ and $x\in U$, there exists some $F\subseteq_{f}X$ such that $x\in(\uparrow F)^{\circ}=\uparrow F\subseteq U$.
Since $G\ll_{n}H$ iff $H\subseteq(\uparrow G)^{\circ}$, definitions of $n$-quasicontinuous and locally hypercompact are equivalent. Similarly, we can define $n$-quasialgebraic spaces. And the following statements hold.
Theorem 5.38.
[9, 19] For a $T_{0}$ topological space $X$, the following are equivalent:
  1. (1)
    $X$ is $n$-quasicontinuous;
  2. (2)
    $X$ is a locally hypercompact space;
  3. (3)
    $\mathcal{O}(X)$ is hypercontinuous.
Definition 5.39.
A $T_{0}$ topological space is called quasicontinuous if one of the four equivalent conditions in Theorem 5.38 holds; called quasialgebraic if it is $n$-quasialgebraic.
Next, we show that for a monotone determined space $X$, $\mathcal{O}(X)$ is algebraic iff it is hyperalgebraic.
Lemma 5.40.
[4, Lemma 2.3] Let $X$ be a topological space and $U,V\in\mathcal{O}(X)$. $U\prec V$ iff $U\subseteq\uparrow F\subseteq V$ for some $F\subseteq_{f}X$.
Theorem 5.41.
[8, Proposition 7] Let $X$ be a monotone determined space. $\mathcal{O}(X)$ is algebraic iff it is hyperalgebraic.
Remark 5.42.
Lemma 5.41 does not always hold for general $T_{0}$ spaces. A simple example is $\mathbb{N}$, the set of natural numbers, endowed with the cofinite topology. Any open subset of it is compact, but not hypercompact.
By the above discussion, we conclude the relationships between the continuity of a $T_{0}$ space and the lattice of its open subsets as follows.
Theorem 5.43.
Let $X$ be a $T_{0}$ space.
  1. (1)
    $X$ is continuous (algebraic) iff $\mathcal{O}(X)$ is a completely distributive lattice (completely distributive algebraic lattice).
  2. (2)
    $X$ is quasicontinuous iff the lattice $\mathcal{O}(X)$ is a hypercontinuous lattice.
  3. (3)
    $X$ is quasialgebraic iff $\mathcal{O}(X)$ is a hyperalgebraic lattice. Moreover, if $X$ is a monotone determined space, then $X$ is quasialgebraic iff $\mathcal{O}(X)$ is an algebraic lattice.
Given any $T_{0}$ space $X$, denote $C(X)$ the lattice of its closed subsets. We show that the quasicontinuity of $X$ can be characterized by the quasicontinuity of $C(X)$, and vice versa. We first recall some equivalent conditions for continuity of lattices.
Lemma 5.44.
[11, Theorem I-3.16] Let $L$ be a complete lattice. The following statements are equivalent:
  1. (1)
    $L$ is completely distributive;
  2. (2)
    $L^{\rm op}$ is completely distributive;
  3. (3)
    $L$ is continuous and every element is the sup of co-primes.
Lemma 5.45.
[11, 10] The following statements hold.
  1. (1)
    Let $L$ be a distributive continuous (algebraic) lattice, then $L$ is hypercontinuous (hyperalgebraic) iff $L^{\rm op}$ is hypercontinuous (quasialgebraic).
  2. (2)
    A complete lattice $L$ is a quasicontinuous (quasialgebraic) lattice iff $\omega(L)$ is a continuous (algebraic) lattice.
Now, it is time to show one of the main results.
Proposition 5.46.
Let $X$ be a $T_{0}$ space. If $C(X)$ is a quasicontinuous lattice, then $X$ is quasicontinuous.
Proof.
Let $C(X)$ be a quasicontinuous domain. Then for any $x\in X$, there exists some $\{A_{1},\dots,A_{n}\}\ll\ \downarrow x$ in $C(X)$. Given any $a_{i}\in A_{i}\ (1\leq i\leq n)$, we have $\{\downarrow a_{1},\dots,\downarrow a_{n}\}\ll\ \downarrow x$.
Claim 1. $\{a_{1},\dots,a_{n}\}\ll_{n}x$ in $X$.
Given any net $(x_{i})_{I}\to x$, $x\in\overline{\bigcup_{i\in I}\{x_{i}\}}$. Then $\downarrow x\subseteq\ \overline{\bigcup_{i\in I}\{x_{i}\}}$. Let =$\mathcal{F}=\{\bigcup_{i\in F}\!\downarrow x_{i}:F\subseteq_{f}I\}$. Then $\mathcal{F}$ is directed and $\downarrow x\leq\bigvee^{\uparrow}\mathcal{F}=\overline{\bigcup_{i\in I}\{x_{i}\}}$ in $C(X)$. There exist some $F\subseteq_{f}I$ and $a_{j}$ such that $\downarrow a_{j}\subseteq\bigcup_{i\in F}\downarrow x_{i}$. Thus, there exists a $i\in F$ such that $a_{j}\leq x_{i}$, i.e., $\{a_{1},\dots,a_{n}\}\ll_{n}x$ in $X$.
Claim 2. For any $x$ and any open neighbourhood $U$ of $x$, there exists some $F\subseteq_{f}U$ such that $x\in(\uparrow F)^{\circ}$.
Given any directed family $\mathcal{A}=\{A_{i}\}_{I}$ of closed subsets of $X$ such that $\downarrow\!x\leq\bigvee\mathcal{A}=\overline{\bigcup_{i\in I}A_{i}}$, then $U\cap(\bigcup_{i\in I}A_{i})\not=\emptyset$. There exists some $i\in I$ such that $A_{i}\cap U\not=\emptyset$. Pick $b\in A_{i}\cap U$, then $\downarrow\!b\subseteq A_{i}$. We have $\{\downarrow\!b:b\in U\}\ll\ \downarrow\!x$ in $C(X)$. Since $C(X)$ is quasicontinuous, there exists a $\{B_{1},\dots,B_{m}\}\subseteq C(X)$ such that $\{\downarrow b:b\in U\}\ll\{B_{1},\dots,B_{m}\}\ll\ \downarrow x$ in $C(X)$. Then
$\forall 1\leq i\leq m,\ \{\downarrow b:b\in U\}\ll B_{i}=\overline{\bigcup_{F\subseteq_{f}B_{i}}\downarrow F},$
which implies that there exists some $F\subseteq_{f}B_{i}$ and $b_{i}\in U$ such that ${\downarrow}b_{i}\subseteq{\downarrow}F\subseteq B_{i}$. Thus, $b_{i}\in B_{i}\cap U$. It follows that $\{{\downarrow}b_{1},\dots,{\downarrow}b_{m}\}\ll\ {\downarrow}x$ in $C(X)$. Then $\{b_{1},\dots,b_{n}\}\ll_{n}x$ by Claim 1, i.e., $x\in({\uparrow}\{b_{1},\dots,b_{n}\})^{\circ}$. ∎
Theorem 5.47.
Let $X$ be a $T_{0}$ space.
  1. (1)
    $X$ is quasicontinuous iff the lattice $C(X)$ is a quasicontinuous lattice.
  2. (2)
    $X$ is quasialgebraic iff the lattice $C(X)$ is a quasialgebraic lattice.
Proof.
(1) By Proposition 5.46, if $C(X)$ is quasicontinuous, then $X$ is quasicontinuous. Conversely, let $X$ be quasicontinuous. Then $\mathcal{O}(X)$ is hypercontinuous by Theorem 5.43. From Theorem 5.45, we know that $C(X)$ is quasicontinuous.
(2) Let $C(X)$ be a quasialgebraic lattice. Then $\omega(C(X))=\upsilon(\mathcal{O}(X))$ is algebraic by Lemma 5.45 and $\mathcal{O}(X)$ is hypercontinuous by (2) and Theorem 5.43. Then $\sigma(\mathcal{O}(X))=\upsilon(\mathcal{O}(X))$ is algebraic and completely distributive. Thus, $\mathcal{O}(X)$ is algebraic and then hyperalgebraic. Conversely, let $X$ be a quasialgebraic space. Then $\mathcal{O}(X)$ is distributive hyperalgebraic lattice. $C(X)$ is a quasialgebraic lattice by Theorem 5.45. ∎
Corollary 5.48.
Let $X$ be a $T_{0}$ topological space. Then the following conditions are equivalent:
  • (1)
    $\mathcal{O}(X)$ with the upper topology is corecompact;
  • (2)
    $\mathcal{O}(X)$ with the upper topology is an injective space;
  • (3)
    $C(X)$ is a quasicontinuous domain;
  • (4)
    $C(X)$ with Lawson topology is Hausdorff;
  • (5)
    $X$ is a quasicontinuous space.
6 Scott completion
The concept of Scott completion is introduced by Zhang, Shi and Li in [25]. It has played an important role in constructing free dcpo algebras over general dcpos, see [4, 6]. But the notion does not seem to have been investigated much. We will give more examples of topological spaces which do not have Scott completions.
Theorem 6.49.
Let $X$ be a $T_{0}$ space and $(\tilde{X},i)$ be a Scott-completion of $X$, then $\mathcal{O}(X)\cong\mathcal{O}(\tilde{X})$.
Proof.
Let $[Y\rightarrow 2]$ denote the set of continuous functions from $Y$ to the Sierpinski space $2$ with the pointwise order. It is well known that $\mathcal{O}(Y)\cong[Y\rightarrow 2]$. For any $f\in[X\rightarrow 2]$, there exists only one $\bar{f}\in[\tilde{X}\rightarrow 2]$ such that $f=\tilde{f}\circ i$ since $(\tilde{X},i)$ is a Scott-completion of $X$. Hence we obtain a map $\theta\colon[X\rightarrow 2]\rightarrow[\tilde{X}\rightarrow 2]$ as follows
$\theta(f)=\bar{f}.$
Clearly $\theta$ is bijective.
$\theta$ is monotone. For any $f_{1},f_{2}\in[X\rightarrow 2]$ with $f_{1}\leq f_{2}$. We need to show that $\bar{f_{1}}\leq\bar{f_{2}}$. Without loss of generality, we consider $y\in\bar{X}$ with $\bar{f_{1}}(y)=1$, i.e., $y\in\bar{f_{1}}^{-1}(1)$. We claim that $i(X)\cap{\downarrow}y\cap\bar{f_{1}}^{-1}(1)\neq\emptyset$. Otherwise we have the indicator function $\mathbbm{1}_{U}$ of $U=({\downarrow}y)^{c}\cap\bar{f_{1}}^{-1}(1)$ with $\mathbbm{1}_{U}\circ i=\bar{f}_{1}\circ i=f$. It follows that $\mathbbm{1}_{U}=\bar{f}_{1}$. Then $U=\mathbbm{1}_{U}^{-1}(1)=\bar{f}^{-1}_{1}(1)$. This is impossible. Back to $i(X)\cap{\downarrow}y\cap\bar{f_{1}}^{-1}(1)\neq\emptyset$. We have some $x\in X$ with $i(x)\in{\downarrow}y\cap\bar{f_{1}}^{-1}(1)$. It means that $f_{1}(x)=\bar{f}_{1}(i(x))=1$. Since $f_{1}\leq f_{2}$, we obtain that $f_{2}(x)=1$. Then $\bar{f}_{2}(y)\geq\bar{f}_{2}(i(x))=f_{2}(x)=1$.
It is easy to see that $\theta$ reflects order. Hence $\theta$ is an order isomorphism. So we obtain that $\mathcal{O}(X)\cong\mathcal{O}(\tilde{X})$. ∎
Remark 6.50.
The above proof is similar to the argument of the order isomorphism between the open lattice of a topological space and the open lattice of its soberification. The map $i^{-1}\colon\mathcal{O}(\tilde{X})\rightarrow\mathcal{O}(X)$ is an order isomorphic map. Furthermore, $i\colon X\rightarrow\tilde{X}$ is an embedding since $i$ is injective, continuous and almost open.
Corollary 6.51.
  1. (1)
    Every $T_{1}$ space endowed with non-discrete topology does not have a Scott completion.
  2. (2)
    Every sober but non-Scott space does not have a Scott-completion.
Proof.
(1) Assume that there is a $T_{1}$ space $X$ with non-discrete topology such that it has a Scott completion $(\Sigma D,i)$, where $D$ is a dcpo. By Theorem 6.49, $\mathcal{O}(X)\cong\sigma(D)$. It follows that $C(X)\cong\Gamma(D)$. Notice that any singleton set is a minimal element in $C(X)\backslash\{\emptyset\}$. Since every minimal nonempty Scott closed of $D$ must be the form of $\{x\}$ ($x\in\text{Min}D$), then $D={\uparrow}\text{Min}D$. It is easy to see that $D$ has the discrete order. Hence $\Gamma(D)$ is a completely distributive lattice. It follows that $X$ is a c-space. Hence $X$ is a discrete space. Contradiction.
(2) Assume that there is a sober space $X$ that is not a Scott space and it has a Scott-completion $(\tilde{X},i)$. Then $i(X)\subsetneqq\tilde{X}$. Consider the soberification map $r\colon\tilde{X}\rightarrow\tilde{X}^{s}$, then $r\circ i\colon X\rightarrow\tilde{X}^{s}$ is also a soberification map. This is impossible. ∎
The following result is a direct application of Theorem 6.49.
Corollary 6.52.
[17, Theorem 4.1] The category of continuous dcpos with Scott continuous maps is not reflective in the category of quasicontinuous dcpos with Scott continuous maps.
Proof.
Notice that $(2,\leq)$ is a continuous domain. The Sierpinski space is $(2,\leq)$ with its Scott topology. Take any quasicontinuous dcpo $D$ which is not continuous. Assume that there exists a continuous dcpo $\tilde{D}$ with a Scott continuous map $i\colon D\rightarrow\tilde{D}$, such that it satisfies the universal property. By the similar argument in Theorem 6.49, we have $\sigma(D)\cong\sigma(\tilde{D})$. This is impossible. ∎
Proposition 6.53.
Let $X$ be a $T_{0}$ space such that it has a Scott-completion $(\tilde{X},i)$. Then $\eta\colon X\rightarrow\Sigma C(X)$ is continuous.
Proof.
Notice that $j=\lambda x.{\downarrow}x\colon\tilde{X}\rightarrow\Sigma\Gamma(\tilde{X})$ is continuous since $\tilde{X}$ is a Scott space. $i^{-1}\colon\Gamma(\tilde{X})\rightarrow C(X)$ is an order isomorphism by Remark 6.50. We claim that
$\eta=i^{-1}\circ j\circ i.$
For any $x\in X$, $i^{-1}\circ j\circ i(x)=i^{-1}({\downarrow}i(x))\supseteq{\downarrow}x=\eta(x)$. On other hand, for any $y\in i^{-1}({\downarrow}i(x))$, i.e., $i(y)\leq i(x)$. Since $i\colon X\rightarrow\tilde{X}$ is an embedding by Remark 6.50, $i$ reflects order. It follows that $y\leq x$. Therefore $\eta$ is continuous. ∎
Corollary 6.54.
Let $X$ be a $T_{0}$ topological space such that its D-completion is a Scott space. Then $\eta\colon X\rightarrow\Sigma C(X)$ is continuous.
7 Open questions
  1. (1)
    Is there a Scott completion $(\tilde{X},i)$ of a topological space $X$, but $(\tilde{X},i)$ is not a D-completion of $X$?
  2. (2)
    Under the conditions of Theorem 4.20, does $\Sigma C(X)\times\Sigma C(X)=\Sigma(C(X)\times C(X))=\Sigma C(X\sqcup X)$ hold?
  3. (3)
    Let $X$ be a $T_{0}$ space and $\Sigma C(X)$ be core-compact. Is $X$ core-compact?
Acknowledgements We thank Yuxu Chen and Peixian Zhou for helpful discussions. We also would like to thank one anonymous referee who points out that one of our main results is not new.

### v2

@@
 Abstract
-We present several equivalent conditions of the continuity of the supremum function $\Sigma C(X)\times\Sigma C(X)\rightarrow\Sigma C(X)$ under mild assumptions, where $C(X)$ denotes the lattice of closed subsets of a $T_{0}$ topological space.
-We also provide an example of a non-monotone determined space $X$ such that $\eta=\lambda x.{\downarrow}x\colon X\rightarrow\Sigma C(X)$ is continuous.
-Additionally, we show that a $T_{0}$ space is quasicontinuous (quasialgebraic) iff the lattice of its closed subsets is a quasicontinuous (quasialgebraic) domain by using $n$-approximation.
+We present several equivalent conditions of the continuity of the supremum function from the square of the Scott space of $C(X)$ to itself under mild assumptions, where $C(X)$ denotes the lattice of closed subsets of a $\mathbf{T_{0}}$ topological space.
+We also show that a $\mathbf{T_{0}}$ space is quasicontinuous (quasialgebraic) iff the lattice of its closed subsets is a quasicontinuous (quasialgebraic) domain by using $n$-approximation.
 Furthermore, we provide a necessary condition for when a topological space possesses a Scott completion.
@@
 1 Introduction
-Topological semilattice has a deep connection with the theory of continuous lattices in domain theory.
-For instance, Lawson showed that compact unital semilattices with small semilattices are exactly continuous lattices with Lawson topology, see [11, Theorem VI-3.4].
-From the viewpoint of category topology, Banaschewski introduced the notion of an essential embedding in the category of $T_{0}$ spaces and continuous maps [2], which has a close relation to Scott’s injective spaces.
-He also obtained a characterization of so-called essentially complete $T_{0}$ spaces in terms of join filters.
-Later Hoffmann provided a more transparent characterization: essentially complete $T_{0}$ spaces are sober spaces which are complete lattices in their specialisation order and satisfy an additional condition connecting the sup-operator with the topology [14].
-In this paper we study a special kind of essentially complete $T_{0}$ space, namely the Scott space of the lattice of closed subsets of a topological space with the supremum operator.
-Directed spaces [24] are special $T_{0}$ topological spaces, extending the concept of Scott spaces.
-The idea is that the priori is a $T_{0}$ space and its convergent directed subsets relative to the specialization order rather than a poset and its existing directed supremums.
-In $T_{0}$ topological spaces, the notion of directed spaces is equivalent to that of monotone determined spaces defined by Erné [8], also is equivalent to crp-spaces introduced by Battenfeld [3, Definition 6.7], in this paper we always assume that topological space is $T_{0}$.
-The category of directed spaces was shown to be Cartesian closed [24, 20].
-$T_{0}$ c-spaces (or $n$-continuous spaces) are all monotone determined spaces.
+Topological semilattice are closely connected to the theory of continuous lattices in domain theory.
+For instance, Lawson showed that compact unital semilattices with small semilattices are exactly continuous lattices equipped with the Lawson topology [11, Theorem VI-3.4].
+From the viewpoint of category topology, Banaschewski introduced the notion of an essential embedding within the category of $\mathbf{T_{0}}$ spaces and continuous maps [2], which has a close relation to Scott’s injective spaces.
+He also obtained a characterization of so-called essentially complete $\mathbf{T_{0}}$ spaces in terms of join filters.
+Subsequently, Hoffmann provided a more transparent characterization: essentially complete $\mathbf{T_{0}}$ spaces are precisely the sober spaces that form complete lattices under their specialization order, satisfying an additional condition relating the supremum operator to the topology [14].
+Powerdomains play an important role in modeling the semantics of nondeterministic programming languages (see [1, Chapter 6] and [11, IV-8]).
+The Hoare powerdomain over a topological space can be realized as the set of its nonempty closed subsets ordered by inclusion.
+In this paper we study a special kind of essentially complete $\mathbf{T_{0}}$ space, namely the Scott space of the lattice of closed subsets of a topological space with the supremum operator.
+Directed spaces [24] constitute a special class of $\mathbf{T_{0}}$ topological spaces which extend the concept of Scott spaces.
+Within $\mathbf{T_{0}}$ spaces, the notion of directed spaces is equivalent both to monotone determined spaces (as established by Erné [8]) and to CRP-spaces (introduced by Battenfeld [3, Definition 6.7]).
+$\mathbf{T_{0}}$ c-spaces (or $n$-continuous spaces) are all monotone determined spaces.
 Based on this, Feng and Kou [9] introduced another approximation relation on monotone determined spaces, called $d$-approximation, and defined the quasicontinuous monotone determined spaces.
 They established an equivalence between quasicontinuous spaces and locally hypercompact spaces.
-Recently, Zhang, Shi and Li [25] have also contributed to this field by introducing the notions of continuous spaces and quasicontinuous spaces, focusing on these spaces within weak monotone convergence spaces.
-It is well known that the hull-kernel topology of the spectrum ${\mathrm{Spec}L}$ equals to the Scott topology when $L$ is a completely distributive (or distributive hypercontinuous) lattice, see [11].
-There is a long-standing open problem asked by Lawson and Mislove that for which distributive continuous lattice its spectrum with the hull-kernel topology is exactly a sober locally compact Scott space (see [23, Problem 528]).
-Chen, Kou and Lyu [4] found a necessary condition for this problem, which is $\sigma(L^{op})=v(L^{op})$.
-Since $L^{op}$ can be viewed as a lattice of closed subsets of a topological space, we continue to investigate the equivalent conditions of $\sigma(C(X))=v(C(X))$ for general topological spaces $X$.
-Surprisingly the continuity of the supremum function $\Sigma C(X)\times\Sigma C(X)\rightarrow\Sigma C(X)$ is one of these equivalent conditions under mild assumptions.
-For monotone determined spaces, we can drop the extra assumption.
+In domain theory, there is a close relationship between the continuity of a dcpo $L$ and that of the lattice $\sigma(L)$ of its Scott-open subsets.
+More surprisingly, the continuity of a dcpo $L$ can be characterized by the continuity of the lattice $\Gamma(L)$ of its Scott-closed subsets (see [4, Theorem 3.17]).
+Moreover, this correspondence extends to general $\mathbf{T_{0}}$ spaces.
+We show that a $\mathbf{T_{0}}$ space $X$ is quasicontinuous (quasialgebraic) iff the lattice of its closed subsets $C(X)$ is quasicontinuous (quasialgebraic) by using $n$-approximation.
+It is well known that for completely distributive (resp., distributive hypercontinuous) lattices $L$, the hull-kernel topology on the spectrum $\mathrm{Spec}L$ coincides with the Scott topology [11].
+A longstanding open problem posed by Lawson and Mislove asks: for which distributive continuous lattices does the spectrum, equipped with the hull-kernel topology, form precisely a sober locally compact Scott space?
+(See [23, Problem 528].) Chen, Kou, and Lyu [4] established a necessary condition for this problem: $\sigma(L^{\mathrm{op}})=v(L^{\mathrm{op}})$.
+Since $L^{\mathrm{op}}$ corresponds to the lattice of closed subsets of a topological space, we investigate equivalent conditions for $\sigma(C(X))=v(C(X))$ in general topological spaces $X$.
+Surprisingly, the continuity of the supremum map $\sup\colon\Sigma C(X)\times\Sigma C(X)\to\Sigma C(X)$ constitutes one such equivalent condition under mild assumptions.
+For monotone determined spaces, the additional assumption can be dropped.
 In addition, we provide an example of non-monotone determined space $X$ such that $\eta=\lambda x.{\downarrow}x\colon X\rightarrow\Sigma C(X)$ is continuous.
-In domain theory, there is a close correspondence between the continuity of a dcpo $L$ and the continuity of the lattice of its Scott open subsets $\sigma(L)$.
-More surprisingly, the continuity of a dcpo $L$ can be characterized by the continuity of its lattice of its Scott closed subsets $\Gamma(L)$ (for example, see [4, Theorem 3.17]).
-These correspondences can be extended to general $T_{0}$ spaces.
-We show that a $T_{0}$ space $X$ is quasicontinuous (quasialgebraic) iff the lattice of its closed subsets $C(X)$ is quasicontinuous (quasialgebraic).
-Monotone convergence spaces are very important in domain theory.
-Every $T_{0}$ space has a free monotone convergence completion called D-completion (for example, see [16]).
+Monotone convergence spaces play a fundamental role in domain theory.
+Every $\mathbf{T_{0}}$ space admits a free monotone convergence completion, known as its D-completion [16].
 We show that D-completion of any monotone determined space is a Scott space.
-The concept of Scott completion is introduced by Zhang, Shi and Li in [25], which has a slight difference with D-completion.
-We also obtain that Scott-completion of any monotone determined space coincides with its D-completion as a corollary.
-The Scott-completion of monotone determined spaces are shown to be very useful to construct free dcpo algebras over any general dcpo (see [6, 7]).
+The concept of Scott completion, introduced by Zhang, Shi, and Li [25], refers to a construction that slightly differs from D-completion.
+As a corollary, we establish that for any monotone determined space, its Scott completion coincides with its D-completion.
+The Scott completion of monotone determined spaces are shown to be very useful to construct free dcpo algebras over any general dcpo (see [6, 7]).
 But the notion does not seem to have been studied much.
@@
 In Section 2, we recall some useful knowledge about monotone determined spaces.
-We devote Section 3 to a discussion of D-completion of monotone determined spaces.
+We discuss D-completion of monotone determined spaces and provide more examples of topological spaces which do not have Scott completions are presented in Section 3.
 In Section 4 the Scott space of the lattice of closed subsets of a topological space with the jointly continuous supremum operator is discussed.
 We characterize the quasicontinuity of lattice of closed subsets in Section 5.
-Finally more examples of topological spaces which do not have Scott completions are presented in Section 6.
 2 Preliminaries
@@
 A nonempty set $P$ endowed with a partial order $\leq$ is called a poset.
-For $A\subseteq P$, we set $\downarrow A=\{x\in P:\exists a\in A,\ x\leq a\}$$\uparrow A=\{x\in P:\exists a\in P,\ a\leq x\}$$A$ is called a lower or an upper set, if $A=\downarrow A$ or $A=\uparrow A$ respectively.
-For an element $a\in P$, we use $\downarrow a$ or $\uparrow a$ instead of $\downarrow\{a\}$ or $\uparrow\{a\}$, respectively.
+For $A\subseteq P$, we set $\downarrow\hskip-2.0ptA=\{x\in P:\exists a\in A,\ x\leq a\}$$\uparrow\hskip-2.0ptA=\{x\in P:\exists a\in P,\ a\leq x\}$AA is called a lower or an upper set, if $A=\downarrow\hskip-2.0ptA$ or $A=\uparrow\hskip-2.0ptA$ respectively.
+For an element $a\in P$, we use $\downarrow\hskip-2.0pta$ or $\uparrow\hskip-2.0pta$ instead of $\downarrow\hskip-2.0pt\{a\}$ or $\uparrow\hskip-2.0pt\{a\}$, respectively.
 Let $P$ be a poset.
 We denote $\sigma(P)$, $\upsilon(P)$ and $A(P)$ to be the Scott topology, the upper topology and the Alexandroff topology on $P$ respectively.
-For a complete lattice $L$, and $x,y\in L$, we say that $x\prec y$ iff $y\in\mathrm{int}_{\upsilon(L)}\uparrow x$.
+For a complete lattice $L$, and $x,y\in L$, we say that $x\prec y$ iff $y\in\mathrm{int}_{\upsilon(L)}\uparrow\hskip-2.0ptx$.
 $L$ is hypercontinuous if $\{d\in L:d\prec x\}$ is directed and has $x$ as a supremum for all $x\in L$.
-$L$ is hyperalgebraic if $x=\bigvee\{d\in L,d\prec d\leq x\}$ for all $x\in L$.
+$L$ is hyperalgebraic if $x=\bigvee\{d\in L:d\prec d\leq x\}$ for all $x\in L$.
 An element $y$ such that $y\prec y$ is called hypercompact.
@@
 For any two topological spaces $X,Y$, we denote $Y^{X}$ or $TOP(X,Y)$ the set of all continuous maps from $X$ to $Y$, endowed with the pointwise order.
-Denote $[X\to Y]_{p}$ and $[X\to Y]_{I}$ to be the topological space equipped with the topology of pointwise convergence and the Isbell topology on $Y^{X}$ respectively.
+Denote $X\to Y]_{p}$ and $X\to Y]_{I}$ to be the topological space equipped with the topology of pointwise convergence and the Isbell topology on $Y^{X}$ respectively.
 We now introduce the notion of a monotone determined space.
-Let $(X,\mathcal{O}(X))$ be a $T_{0}$ space.
+Let $(X,\mathcal{O}(X))$ be a $\mathbf{T_{0}}$ space.
 Every directed subset $D\subseteq X$ can be regarded as a monotone net $(d)_{d\in D}$.
@@
 Lemma 2.1.
-Let $X$ be a $T_{0}$ space. For any $(D,x)\in DS(X)\times X$, $D\rightarrow x$ if and only if $D\cap U\not=\emptyset$ for any open neighborhood of $x$.
-Set $DLim(X)=\{(D,x)\in DS(X)\times X:\ D\rightarrow x\}$ to be the set of all pairs of directed subsets and their limits in $X$. Then ($(\{y\},x)\in DLim(X)$ iff $x\sqsubseteq y$ for all $x,y\in X$.
+Let $X$ be a $\mathbf{T_{0}}$ space. For any $(D,x)\in DS(X)\times X$, $D\rightarrow x$ if and only if $D\cap U\not$$\emptyset$ for any open neighborhood of $x$.
+Set $DLim(X)=\{(D,x)\in DS(X)\times X:\ D\rightarrow x\}$ to be the set of all pairs of directed subsets and their limits in $X$. Then $(\{y\},x)\in DLim(X)$ iff $x\sqsubseteq y$ for all $x,y\in X$.
 Definition 2.2.
-Let $X$ be a $T_{0}$ space. A subset $U\subseteq X$ is called directed-open if for all $(D,x)\in DLim(X)$, $x\in U$ implies $D\cap U\not=\emptyset$.
+Let $X$ be a $\mathbf{T_{0}}$ space. A subset $U\subseteq X$ is called directed-open if for all $(D,x)\in DLim(X)$, $x\in U$ implies $D\cap U\not$$\emptyset$.
 Obviously, every open set of $X$ is directed-open. Set $d(\mathcal{O}(X))=\{U\subseteq X:U\ {\rm is}\ {\rm directed}\text{-}{\rm open}\},$ then $\mathcal{O}(X)\subseteq d(\mathcal{O}(X))$.
 Theorem 2.3.
-[24] Let $X$ be a $T_{0}$ topological space. Then
+[24, Theorem 2.4] Let $X$ be a $\mathbf{T_{0}}$ topological space. Then
   1. (1)
-    For all $U\in d(\mathcal{O}(X))$, $U=\uparrow U$;
+    For all $\in d(\mathcal{O}(X)), U=U$$\uparrow\hskip-2.0ptU$;
   2. (2)
-    $X$ equipped with $d(\mathcal{O}(X))$ is a $T_{0}$ topological space such that $\sqsubseteq_{d}=\sqsubseteq$, where $\sqsubseteq_{d}$ is the specialization order relative to $d(\mathcal{O}(X))$.
+    $X$ equipped with $d(\mathcal{O}(X))$ is a $\mathbf{T_{0}}$ topological space such that $\sqsubseteq_{d}=\sqsubseteq$, where $\sqsubseteq_{d}$ is the specialization order relative to $d(\mathcal{O}(X))$.
   3. (3)
@@
 Definition 2.4.
-[24] A topological space $X$ is said to be a monotone determined space if it is $T_{0}$ and every directed-open set is open; equivalently, $d(\mathcal{O}(X))=\mathcal{O}(X)$.
+[24, Definition 3.1] A topological space $X$ is said to be a monotone determined space if it is $\mathbf{T_{0}}$ and every directed-open set is open; equivalently, $d(\mathcal{O}(X))=\mathcal{O}(X)$.
 Given any space $X$, we denote $\mathcal{D}X$ the topological space $(X,d(\mathcal{O}(X)))$.
 Theorem 2.5.
-[24] Let $X$ be a $T_{0}$ space, we have
+[24, Theorem 3.2] Let $X$ be a $\mathbf{T_{0}}$ space, we have
   1. (1)
@@
     2. (ii)
-      For all $U\subseteq X$, $U$ is open iff for any $(D,x)\in DLim(X)$, $x\in U$ implies $U\cap D\not=\emptyset$.
+      For all $\subseteq X, UU$ is open iff for any $(D,x)\in DLim(X)$, $x\in U$ implies $U\cap D\not$$\emptyset$.
     3. (iii)
-      For all $A\subseteq X$, $A$ is closed iff for any directed subset $D\subseteq A$, $D\rightarrow x$ implies $x\in A$ for all $x\in X$.
+      For all $\subseteq X, AA$ is closed iff for any directed subset $D\subseteq A$, $D\rightarrow x$ implies $x\in A$ for all $x\in X$.
 Monotone determined spaces include many important examples in domain theory.
@@
   2. (2)
-    ([8]) Every poset endowed with the weak Scott topology is a monotone determined space.
+    ([8, Section 7]) Every poset endowed with the weak Scott topology is a monotone determined space.
   3. (3)
@@
 We denote $\bf DTop$} the category of all nonempty monotone determined spaces with continuous maps as morphisms.
-Lemma 2.7.
-Given any two monotone determined spaces $X,Y$, a subset $U$ of $X\otimes Y$ is open iff the following conditions hold:
-  1. (1)
-    for any directed subsets $(x_{i})_{I}$ with $(x_{i})_{I}\to x$ in $X$ and any $(x,y)\in U$, we have $((x_{i},y))_{I}\cap U\neq\emptyset$;
-  2. (2)
-    for any directed subsets $(y_{i})_{I}$ with $(y_{i})_{I}\to y$ in $Y$ and any $(x,y)\in U$ we have $((x,y_{i}))_{I}\cap U\neq\emptyset$.
-Proof.
-Assume that $U$ is open in $X\otimes Y$ and $(x,y)\in U$.
-For any directed set $(x_{i})_{I}\to x$ in $X$ and any $y\in Y$, $((x_{i},y))_{I}$ is a directed subset of $X\otimes Y$ and $((x_{i},y))_{I}\to(x,y)$ in $X\times Y$.
-By Theorem 2.3 (3), $((x_{i},y))_{I}\to(x,y)$ in $X\otimes Y$.
-Thus, $((x_{i},y))_{I}\cap U\neq\emptyset$.
-It is the same for $(y_{i})_{I}$.
-Conversely, assume that $(1)$ and $(2)$ are satisfied.
-We show that $U$ is open in $X\otimes Y$.
-It is easily seen that $U$ is an upper set relative to the specialization order of $X\times Y$.
-Let $D=((x_{i},y_{i}))_{I}$ be a directed subset of $X\times Y$ and converges to $(x,y)\in U$ in $X\times Y$.
-We have that $(x_{i})_{I}\to x$ in $X$ and $(y_{i})_{I}\to y$ in $Y$ respectively.
-Thus, $((x_{i},y))_{I}\to(x,y)$ in $X\times Y$, and then there exists some $i_{0}\in I$ such that $(x_{i_{0}},y)\in U$.
-By $((x_{i_{0}},y_{i}))_{I}\to(x_{i_{0}},y)$, there exists some $i_{1}\in I$ such that $(x_{i_{0}},y_{i_{1}})\in U$.
-Let $i_{0},i_{1}\leq i_{2}$, then $(x_{i_{2}},y_{i_{2}})\in U$.
-$\Box$
-3 D-completion of monotone determined spaces
+3 D-completion and Scott completion
 Monotone convergence spaces (a.k.a d-spaces) is a very important class of topological spaces in domain theory.
-It is well known that every $T_{0}$ space has a D-completion.
+It is well known that every $\mathbf{T_{0}}$ space has a D-completion.
 Let us recall some results about D-completion for further discussions.
 Definition 3.7.
-Let $X$ be a $T_{0}$ space. If $X$ with the specialization order is a dcpo and every open set of $X$ is Scott open in $(X,\leq)$, then we call $X$ a monotone convergence space space.
+Let $X$ be a $\mathbf{T_{0}}$ space. If $X$ with the specialization order is a dcpo and every open set of $X$ is Scott open in $(X,\leq)$, then we call $X$ a monotone convergence space space.
 A subset $A$ of a poset $P$ is called d-closed if for every directed set $D$ of $A$ that possesses supremum $\bigvee D$, $\bigvee D$ is included in $A$.
-The d-closed sets form the closed sets for a topology, called the d-topology [16].
+The d-closed sets form the closed sets for a topology, called the d-topology [16, Section 5].
 We denote the closure of $A$ in $P$ with d-topology by $\text{cl}_{d}(A)$.
@@
 Definition 3.8.
-[16] An embedding $j\colon X\rightarrow\tilde{X}$ of a space $X$ with image a d-dense subset of a d-space $\tilde{X}$ is called a $D$-completion.
+[16, Definition 6.5] An embedding $j\colon X\rightarrow\tilde{X}$ of a space $X$ with image a d-dense subset of a d-space $\tilde{X}$ is called a $D$-completion.
 Lemma 3.9.
-[16] Consider the following properties for a subset $A$ of a $T_{0}$ space
+[16, Lemma 6.3] Consider the following properties for a subset $A$ of a $\mathbf{T_{0}}$ space
   1. (1)
-    $A$ is a monotone convergence space.
+    AA is a monotone convergence space.
   2. (2)
-    $A$ is a sub-dcpo.
+    AA is a sub-dcpo.
   3. (3)
-    $A$ is d-closed.
+    AA is d-closed.
 Then $(1)$ implies $(2)$ implies $(3)$, and all three are equivalent if $X$ is a monotone convergence space.
@@
 Theorem 3.10.
-[16] Let $j\colon X\rightarrow Y$ be a topological embedding of a space $X$ into a monotone convergence space $Y$.
+[16, Theorem 6.7] Let $j\colon X\rightarrow Y$ be a topological embedding of a space $X$ into a monotone convergence space $Y$.
 Let $\tilde{X}=\text{cl}_{d}(j(X))$ be the d-closure of $j(X)$ in $Y$, equipped with the relative topology from $Y$.
@@
 Remark 3.11.
-For any $T_{0}$ space $X$, there is a natural monotone convergence space $C(X)$ (the lattice of closed subsets of $X$) with the upper topology such that $\eta=\lambda x.{\downarrow}x\colon X\rightarrow C(X)$ is an embedding.
+For any $\mathbf{T_{0}}$ space $X$, there is a natural monotone convergence space $C(X)$ (the lattice of closed subsets of $X$) with the upper topology such that $\eta=\lambda x.{\downarrow}x\colon X\rightarrow C(X)$ is an embedding.
 We call the d-closure of $\eta(X)$ in $C(X)$ the standard D-completion of $X$ and denote it by $X^{d}$.
 Proposition 3.12.
-Let $X$ be a $T_{0}$ space and $X^{d}$ be its D-completion. Then $\mathcal{O}(X)\cong\mathcal{O}(X^{d})$.
+Let $X$ be a $\mathbf{T_{0}}$ space and $X^{d}$ be its D-completion. Then $\mathcal{O}(X)\cong\mathcal{O}(X^{d})$.
 Proof.
@@
 For any Scott open set $\mathcal{U}$ of $(X^{d},\leq)$ and any closed set $A\in\mathcal{U}$, notice that
-[…]
+$A\in X^{d}=\text{cl}_{d}\{{\downarrow}x:x\in X\}=\text{cl}_{d}\{{\downarrow}x:x\in A\}\cup\text{cl}_{d}\{{\downarrow}x:x\in X\backslash A\}.$
 […]
@@
 Example 3.16.
-Let $D$ be a set ×$\mathbb{N}\times\mathbb{N}\cup\{a_{n}:n\in\mathbb{N}\}\cup\{\top\}$. The order on $D$ is defined as follows: $x\leq y$ iff
+Let $D$ be a set $\mathbb{N}\times\mathbb{N}\cup\{a_{n}:n\in\mathbb{N}\}\cup\{\top\}$. The order on $D$ is defined as follows: $x\leq y$ iff
   • 1.
@@
     There are natural numbers $n$ and $m$ such that $x=(n,m),y=a_{n}$.
+$D$ can be easily depicted as in Figure 1.
+Let $X=\mathbb{N}\times\mathbb{N}\cup\{\top\}$ with the subspace topology of $\Sigma D$.
+It is easy to see that $\Sigma D$ is a D-completion of $X$.
+Notice that $\{\top\}$ is always a directed open set of $X$ which is not open in $X$.
+Hence $X$ is not a monotone determined space.
+$a_{1}$$a_{2}$$a_{3}$$a_{3}\top$
+Figure 1: $\mathbb{N}\times\mathbb{N}\cup\{a_{n}:n\in\mathbb{N}\}\cup\{\top\}$
+The concept of Scott completion is introduced by Zhang, Shi and Li in [25].
+It has played an important role in constructing free dcpo algebras over general dcpos, see [4, 6].
+But the notion does not seem to have been investigated much.
+We will give more examples of topological spaces which do not have Scott completions in this section.
+Definition 3.17.
+[25] A Scott completion $(Y,f)$ of a space $X$ is a Scott space $Y$ together with a continuous map $f\colon X\rightarrow Y$ such that for any Scott space $Z$ and continuous map $g\colon X\rightarrow Z$, there exists a unique continuous map $\tilde{g}$ satisfying $g=\tilde{g}\circ f$.
+Theorem 3.18.
 […]
-$a_{1}$$a_{2}$$a_{3}$$a_{3}$$\top$
-Figure 1: ×$\mathbb{N}\times\mathbb{N}\cup\{a_{n}:n\in\mathbb{N}\}\cup\{\top\}$
+Proof.
+Let $Y\rightarrow 2]$ denote the set of continuous functions from $Y$ to the Sierpinski space $2$ with the pointwise order.
+It is well known that $\mathcal{O}(Y)\cong[Y\rightarrow 2]$.
+For any $f\in[X\rightarrow 2]$, there exists only one $\bar{f}\in[\tilde{X}\rightarrow 2]$ such that $f=\tilde{f}\circ i$ since $(\tilde{X},i)$ is a Scott-completion of $X$.
+Hence we obtain a map $\theta\colon[X\rightarrow 2]\rightarrow[\tilde{X}\rightarrow 2]$ as follows
+$\theta(f)=\bar{f}.$
+Clearly $\theta$ is bijective.
+$\theta$ is monotone.
+For any $f_{1},f_{2}\in[X\rightarrow 2]$ with $f_{1}\leq f_{2}$.
+We need to show that $\bar{f_{1}}\leq\bar{f_{2}}$.
+Without loss of generality, we consider $y\in\bar{X}$ with $\bar{f_{1}}(y)=1$, i.e., $y\in\bar{f_{1}}^{-1}(1)$.
+We claim that $i(X)\cap{\downarrow}y\cap\bar{f_{1}}^{-1}(1)\neq\emptyset$.
+Otherwise we have the indicator function $\mathbbm{1}_{U}$ of $U=({\downarrow}y)^{c}\cap\bar{f_{1}}^{-1}(1)$ with $\mathbbm{1}_{U}\circ i=\bar{f}_{1}\circ i=f$.
+It follows that $\mathbbm{1}_{U}=\bar{f}_{1}$.
+Then $U=\mathbbm{1}_{U}^{-1}(1)=\bar{f}^{-1}_{1}(1)$.
+This is impossible.
+Back to $i(X)\cap{\downarrow}y\cap\bar{f_{1}}^{-1}(1)\neq\emptyset$.
+We have some $x\in X$ with $i(x)\in{\downarrow}y\cap\bar{f_{1}}^{-1}(1)$.
+It means that $f_{1}(x)=\bar{f}_{1}(i(x))=1$.
+Since $f_{1}\leq f_{2}$, we obtain that $f_{2}(x)=1$.
+Then $\bar{f}_{2}(y)\geq\bar{f}_{2}(i(x))=f_{2}(x)=1$.
+It is easy to see that $\theta$ reflects order. Hence $\theta$ is an order isomorphism. So we obtain that $𝒪(X~)\mathcal$$\cong\mathcal{O}(\tilde{X})$. ∎
+Remark 3.19.
+The above proof is similar to the argument of the order isomorphism between the open lattice of a topological space and the open lattice of its soberification.
+The map $i^{-1}\colon\mathcal{O}(\tilde{X})\rightarrow\mathcal{O}(X)$ is an order isomorphic map.
+Furthermore, $i\colon X\rightarrow\tilde{X}$ is an embedding since $i$ is injective, continuous and almost open.
+Corollary 3.20.
+  1. (1)
+    Every $T_{1}$ space endowed with non-discrete topology does not have a Scott completion.
+  2. (2)
+    Every sober but non-Scott space does not have a Scott-completion.
+Proof.
+[…]
+(2) Assume that there is a sober space $X$ that is not a Scott space and it has a Scott-completion $(\tilde{X},i)$.
+Then $i(X)\subsetneqq\tilde{X}$.
+Consider the soberification map $r\colon\tilde{X}\rightarrow\tilde{X}^{s}$, then $r\circ i\colon X\rightarrow\tilde{X}^{s}$ is also a soberification map.
+This is impossible.
+∎
+The following result is a direct application of Theorem 3.18.
+Corollary 3.21.
+[17, Theorem 4.1] The category of continuous dcpos with Scott continuous maps is not reflective in the category of quasicontinuous dcpos with Scott continuous maps.
+Proof.
+[…]
+Proposition 3.22.
+Let $X$ be a $\mathbf{T_{0}}$ space such that it has a Scott-completion $(\tilde{X},i)$. Then $\eta\colon X\rightarrow\Sigma C(X)$ is continuous.
+Proof.
+[…]
+$\eta=i^{-1}\circ j\circ i.$
+For any $x\in X$, $i^{-1}\circ j\circ i(x)=i^{-1}({\downarrow}i(x))\supseteq{\downarrow}x=\eta(x)$.
+On other hand, for any $y\in i^{-1}({\downarrow}i(x))$, i.e., $i(y)\leq i(x)$.
+Since $i\colon X\rightarrow\tilde{X}$ is an embedding by Remark 3.19, $i$ reflects order.
+It follows that $y\leq x$.
+Therefore $\eta$ is continuous.
+∎
+Corollary 3.23.
+Let $X$ be a $\mathbf{T_{0}}$ topological space such that its D-completion is a Scott space. Then $\eta\colon X\rightarrow\Sigma C(X)$ is continuous.
+Example 3.24.
+There exists a topological space $X$ such that $\eta\colon X\rightarrow\Sigma C(X)$ is continuous while $X$ does not have a Scott completion.
+Let $\mathcal{J}$ be Johnstone’s non-sober dcpo [12, Exercise 5.2.15].
+Notice that $Id(\prod^{n}\mathcal{J})=\prod^{n}Id(\mathcal{J})$}, by using Corollary 4.32 and Proposition 4.35 we obtain that $\sigma(\Gamma(\mathcal{J}))=v(\Gamma(\mathcal{J}))$.
+Now let $X$ be the soberification of $\Sigma\mathcal{J}$.
+It implies that $\sigma(C(X))=v(C(X))$.
+Since $X$ is not a Scott space, $X$ does not have a Scott completion by Corollary 3.20.
+It is easy to verify that $\eta\colon X\rightarrow\Sigma C(X)$ is continuous.
 4 Continuity of the supremum function $\Sigma C(X)\times\Sigma C(X)\rightarrow\Sigma C(X)$
@@
 For any $n\in\mathbb{N}$, define a map $s_{n}:\prod\limits^{n}X\to\Sigma(C(X))$ as follows: $\forall(x_{1},x_{2},\dots,x_{n})\in\prod\limits^{n}X$,
-$s_{n}(x_{1},x_{2},\dots,x_{n})=\ \downarrow\!{\{x_{1},x_{2},\dots,x_{n}\}}.$
+$s_{n}(x_{1},x_{2},\dots,x_{n})=\ \downarrow\hskip-2.0pt\!{\{x_{1},x_{2},\dots,x_{n}\}}.$
 The following result is due to R.Hoffmann.
 Theorem 4.25.
-[14, Proposition 1.1] A space is an essentially complete $T_{0}$ space iff the following are satisfied:
+[14, Proposition 1.1] A space is an essentially complete $\mathbf{T_{0}}$ space iff the following are satisfied:
   1. (i)
-    The associated pre-order $(X,\leq)$ is a complete lattice.
+    The associated pre-order (i.e., specialization order) $(X,\leq)$ is a complete lattice.
   2. (ii)
@@
     Whenever $a=\sup\{b,c\}$ with $a,b,c\in X$, then $\{U\cap V|b\in U\in\mathcal{O}(X),c\in U\in\mathcal{O}(X)\}$ is an open neighborhood basis for $a$ in $X$.
-Hoffmann noticed that the above condition $(iii)$ can be replaced by that “$\sup\colon X\times X\rightarrow X$ is continuous” [14, Lemma 1.5].
-Proposition 4.19.
-Let $X,Y,Z$ be monotone determined spaces. A map $f:X\otimes Y\to Z$ is continuous iff it is separately continuous.
+Hoffmann noticed that the above condition $(iii)$ can be replaced by that $\sup\colon X\times X\rightarrow X$ is continuous” [14, Lemma 1.5].
 Theorem 4.26.
 For a topological space $X$, the following conditions are equivalent:
-  • (1)
+  1. (1)
     $\sigma(C(X))=\upsilon(C(X))$;
-  • (2)
+  2. (2)
     $s_{n}$ is continuous for all $n\in\mathbb{N}$;
-  • (3)
+  3. (3)
     $s_{1}$ is continuous and $\bigcup\colon\Sigma C(X)\times\Sigma C(X)\rightarrow\Sigma C(X)$ is continuous.
@@
 Let $\mathcal{U}$ be an open subset of $\Sigma(C(X))$ and $A\in\mathcal{U}$.
-Without loss of generality, we assume $A\not=\emptyset$.
-Note that since $A=\bigcup\{\downarrow\!F:F\subseteq_{f}\!A\}$ and $\{\downarrow F:F\subseteq_{f}A\}$ is a directed family in $C(X)$, there exists a non-empty finite subset $F$ of $A$ such that $\downarrow\!F\in\mathcal{U}$.
-Let $F=\{x_{1},x_{2},\ldots,x_{n}\}$, then $s_{n}(x_{1},x_{2},\ldots,x_{n})=\downarrow\!F\in\mathcal{U}$.
+Without loss of generality, we assume $A\not$$\emptyset$.
+Note that since $A=\bigcup\{\downarrow\hskip-2.0pt\!F:F\subseteq_{f}\!A\}$ and $\{\downarrow\hskip-2.0ptF:F\subseteq_{f}A\}$ is a directed family in $C(X)$, there exists a non-empty finite subset $F$ of $A$ such that $\downarrow\hskip-2.0pt\!F\in\mathcal{U}$.
+Let $F=\{x_{1},x_{2},\ldots,x_{n}\}$, then $s_{n}(x_{1},x_{2},\ldots,x_{n})=\downarrow\hskip-2.0pt\!F\in\mathcal{U}$.
 It follows that $(x_{1},x_{2},\ldots,x_{n})\in s_{n}^{-1}(\mathcal{U})$.
@@
 $(x_{1},x_{2},\ldots,x_{n})\in U_{1}\times U_{2}\times\cdots\times U_{n}\subseteq s_{n}^{-1}(\mathcal{U}).$
-Since $x_{k}\in A$ for $1\leq k\leq n$, we have $A\in\Diamond U_{k}=\{B\in C(X):\ B\cap U_{k}\not=\emptyset\}$.
+Since $x_{k}\in A$ for $1\leq k\leq n$, we have $A\in\Diamond U_{k}=\{B\in C(X):\ B\cap U_{k}\not$$\emptyset\}$.
 It follows that $A\in\bigcap\limits_{k=1}^{n}\Diamond U_{k}\in\upsilon(C(X))$.
 For any $B\in\bigcap\limits_{k=1}^{n}\Diamond U_{k}$, there exists $y_{k}\in B\cap U_{k}$ for $1\leq k\leq n$.
-Since $(y_{1},y_{2},\dots,y_{n})\in s_{n}^{-1}(\mathcal{U})$, we have $\bigcup\limits_{k=1}^{n}\downarrow y_{k}\in\mathcal{U}$.
+Since $(y_{1},y_{2},\dots,y_{n})\in s_{n}^{-1}(\mathcal{U})$, we have $\bigcup\limits_{k=1}^{n}\downarrow\hskip-2.0pty_{k}\in\mathcal{U}$.
 It follows that $B\in\mathcal{U}$, i.e., $A\in\bigcap\limits_{k=1}^{n}\Diamond U_{k}\subseteq\mathcal{U}$.
@@
 The equivalence of $(1)$ and $(2)$ has already been given in [5, Proposition 4.1].
-[…]
+$\top\bot$
 Figure 2: P
@@
 Let $X$ be a monotone determined space. The following statements are equivalent:
-  • (1)
+  1. (1)
     $\sigma(C(X))=\upsilon(C(X))$;
-  • (2)
+  2. (2)
     $s_{n}$ is continuous for all $n\in\mathbb{N}$;
-  • (3)
+  3. (3)
     $\bigcup\colon\Sigma C(X)\times\Sigma C(X)\rightarrow\Sigma C(X)$ is continuous.
@@
 By a similar argument in Theorem 4.26, we can show that $\tau$ is equal to $v(C(X))$ if the operator $\bigcup\colon(C(X),\tau)\times(C(X),\tau)\rightarrow(C(X),\tau)$ is continuous.
-In other words, there exists only one topology $\tau$ on $C(X)$ such that $(1)$ the specialization order of $\tau$ is equal to the inclusion order and $(2)$ $(C(X),\tau)$ is an essentially complete $T_{0}$ space.
+In other words, there exists only one topology $\tau$ on $C(X)$ such that $(1)$ the specialization order of $\tau$ is equal to the inclusion order and $(2)$ $(C(X),\tau)$ is an essentially complete $\mathbf{T_{0}}$ space.
 But this is false for a non-monotone determined space $X$, for instance see Example 4.28.
@@
 Let $D$ be a poset. The following conditions are equivalent:
-  • (1)
+  1. (1)
     $\sigma(\Gamma(D))=\upsilon(\Gamma(D))$;
-  • (2)
+  2. (2)
     $s_{n}$ is continuous for all $n\in\mathbb{N}$;
-  • (3)
+  3. (3)
     $\bigcup\colon\Sigma\Gamma(D)\times\Sigma\Gamma(D)\rightarrow\Sigma\Gamma(D)$ is continuous.
@@
 Corollary 4.32.
-[4, Proposition 3.5] Let $P$ be a poset. If ${\displaystyle\Sigma(\prod^{n}P)=\prod^{n}(\Sigma P)}$ for each $n\in\mathbb{N}$, then $\sigma(\Gamma(P))=v(\Gamma(P))$.
+[4, Proposition 3.5] Let $P$ be a poset. If $\Sigma(\prod^{n}P)=\prod^{n}(\Sigma P)$} for each $n\in\mathbb{N}$, then $\sigma(\Gamma(P))=v(\Gamma(P))$.
 Proposition 4.33.
@@
 Remark 4.34.
-$(2)$ implies $(1)$ can be found in [4, Proposition 3.15].The above results does not hold for $T_{0}$ spaces.
+$(2)$ implies $(1)$ can be found in [4, Proposition 3.15].The above results do not hold for $\mathbf{T_{0}}$ spaces.
 For example we consider $X=\Sigma L$ where $L$ is the Isbell’s lattice, then $(3)$ holds while $(2)$ does not hold since $\Sigma C(X)$ is not sober [21].
@@
 We consider the space $X$ in Example 3.16, it follows that $C(X)\cong\Gamma(D)$ since $\Sigma D$ is the D-completion of $X$.
-Notice that ${\displaystyle Id(\prod^{n}D)=\prod^{n}Id(D)}$, by using Corollary 4.26 and Proposition 4.29 we obtain that $\sigma(\Gamma(D))=v(\Gamma(D))$.
+Notice that $Id(\prod^{n}D)=\prod^{n}Id(D)$}, by using Corollary 4.32 and Proposition 4.35 we obtain that $\sigma(\Gamma(D))=v(\Gamma(D))$.
 Hence we also have $\sigma(C(X))=v(C(X))$.
@@
 By the above lemma, if $\Sigma L\times\Sigma M$ is not a Scott space, then $\bigvee\colon\Sigma S\times\Sigma S\rightarrow\Sigma S$ is not continuous.
-We recall that a topological space $X$ is called a Scott space if its topology is equal to the Scott topology on $(X,\leq)$, which $\leq$ is the specialization order of $X$.
+We recall that a topological space $X$ is called a Scott space if its topology is equal to the Scott topology on $(X,\leq)$, where $\leq$ is the specialization order of $X$.
 Proposition 4.38.
-[11, Proposition II-3.15(iii)] If $X$ is a continuous retract of a monotone convergence space $Y$, then $X$ must be $T_{0}$ and is a monotone convergence space.
+[11, Proposition II-3.15(iii)] If $X$ is a continuous retract of a monotone convergence space $Y$, then $X$ must be $\mathbf{T_{0}}$ and is a monotone convergence space.
 In the special case that $Y$ is a dcpo equipped with the Scott topology, then $X$ must also have the Scott topology.
@@
 We use Hertling’s example [13, Theorem 7.4] to show that there exists a complete lattice $Z$ such that condition $(2)$ holds but condition $(1)$ fails.
-$Z$ is the disjoint union of a complete lattice $X$ and $Y=\sigma(X)$ plus a top element and a bottom element, i.e., $Z=\{X+Y\}\bigcup\{\top,\bot\}$.
-$Z$ has the following conditions:
+$Z$ is the disjoint union of a complete lattice $X$ and $Y=\sigma(X)$ plus a top element and a bottom element, i.e., $\{X+Y\}\bigcup\{\top,\bot\}.
+ZZ$ has the following conditions:
   • 1.
@@
 We introduce the concepts of quasicontinuous spaces and quasialgebraic spaces.
-Then we will summarize the correspondences among the continuity of a $T_{0}$ space, the continuity of the lattice of its open subsets, and the continuity of the lattice of its closed subsets.
+Then we will summarize the correspondences among the continuity of a $\mathbf{T_{0}}$ space, the continuity of the lattice of its open subsets, and the continuity of the lattice of its closed subsets.
 Given any set $X$, we use $F\subseteq_{f}X$ to denote that $F$ is a finite subset of $X$.
-Given any two subsets $G,H\subseteq X$, we define $G\leq H$ iff $\uparrow H\subseteq\uparrow G$.
+Given any two subsets $G,H\subseteq X$, we define $G\leq H$ iff $\uparrow\hskip-2.0ptH\subseteq\uparrow\hskip-2.0ptG$.
 A family of finite sets is said to be directed if given $F_{1},F_{2}$ in the family, there exists an $F$ in the family such that $F_{1},F_{2}\leq F$.
-Let $X$ be a $T_{0}$ topological space and $\mathcal{M}$ be a directed family of finite subsets of $X$.
-Then $F_{\mathcal{M}}=\{A\subseteq X:\exists M\in\mathcal{M},\uparrow M\subseteq A\}$ is a filter under the inclusion order.
+Let $X$ be a $\mathbf{T_{0}}$ topological space and $\mathcal{M}$ be a directed family of finite subsets of $X$.
+Then $F_{\mathcal{M}}=\{A\subseteq X:\exists M\in\mathcal{M},\uparrow\hskip-2.0ptM\subseteq A\}$ is a filter under the inclusion order.
 We say that $\mathcal{M}$ converges to $x$ if $F_{\mathcal{M}}$ converges to $x$, i.e., for any open neighbourhood $U$ of $x$, there exists some $A\in\mathcal{M}$ such that $A\subseteq U$.
 Definition 5.40.
-Let $X$ be a $T_{0}$ space and $G,H\subseteq X$.
-We say that $G$ $n$-$approximates$ $H$, denoted by $G\ll_{n}H$, if for every net $(x_{j})_{J}$ of $X$, $(x_{j})_{J}\to y$ for some $y\in H$ implies $(x_{j})_{J}\cap\uparrow G\neq\emptyset$.
-We write $G\ll_{n}x$ for $G\ll_{n}\{x\}$.
-$G$ is said to be $n$-compact if $G$ is a finite set and $G\ll_{n}G$.
+Let $X$ be a $\mathbf{T_{0}}$ space and $G,H\subseteq X$.
+We say that $G$ $n$-approximatesapproximates $H$, denoted by $G\ll_{n}H$, if for every net $(x_{j})_{J}$ of $X$, $(x_{j})_{J}\to y$ for some $y\in H$ implies $(x_{j})_{J}\cap\uparrow\hskip-2.0ptG\neq\emptyset$.
+We write $G\ll_{n}x$ for $\ll_{n}\{x\}.
+GG$ is said to be $n$-compact if $G$ is a finite set and $G\ll_{n}G$.
 Lemma 5.41.
-For any $T_{0}$ space $X$, the following are equivalent: for all $x,y\in Y$,
+For any $\mathbf{T_{0}}$ space $X$, the following are equivalent: for all $x,y\in Y$,
   1. (1)
@@
   2. (2)
-    $H\subseteq(\uparrow G)^{\circ}$, where $(\uparrow G)^{\circ}$ means the interior of $\uparrow G$.
+    $H\subseteq(\uparrow\hskip-2.0ptG)^{\circ}$, where $(\uparrow\hskip-2.0ptG)^{\circ}$ means the interior of $\uparrow\hskip-2.0ptG$.
 Proof.
@@
 We only show $(1)\Rightarrow(2)$.
-Assume that $y\in H,y\not\in(\uparrow G)^{\circ}$.
-Then, for any open set $U\subseteq X$ with $y\in U$, there exists a $y_{U}\in U$ such that $y_{U}\not\in\uparrow G$.
-Set $J=\{U\in\mathcal{O}(X):y\in U\}$ ordered as: $\forall U,V\in J$,$U\leq V\Leftrightarrow V\subseteq U.$ Then $J$ is directed.
+Assume that $y\in H,y\not\in(\uparrow\hskip-2.0ptG)^{\circ}$.
+Then, for any open set $U\subseteq X$ with $y\in U$, there exists a $y_{U}\in U$ such that $y_{U}\not\in\uparrow\hskip-2.0ptG$.
+Set $J=\{U\in\mathcal{O}(X):y\in U\}$ ordered as: $\forall U,V\in J$$U\leq V\Leftrightarrow V\subseteq U.$ Then $J$ is directed.
 Easily, one can see that $(y_{U})_{U\in J}$ is a net converging to $y$.
 Since $G\ll_{n}y$, there exists $U_{0}\in J$ such that $x\sqsubseteq y_{U_{0}}$ for some $x\in G$ by definition.
-It is a contradiction for $y_{U_{0}}\not\in\uparrow G$.
-Hence, $y\in(\uparrow G)^{\circ}$.
+It is a contradiction for $y_{U_{0}}\not\in\uparrow\hskip-2.0ptG$.
+Hence, $y\in(\uparrow\hskip-2.0ptG)^{\circ}$.
 $\Box$
-Since the $n$-approximation does not depend on monotone determined space, a notion of $n$-quasicontinuity for a $T_{0}$ space can be defined as follows.
+Since the $n$-approximation does not depend on monotone determined space, a notion of $n$-quasicontinuity for a $\mathbf{T_{0}}$ space can be defined as follows.
 Definition 5.42.
-A $T_{0}$ topological space $X$ is called $n$-quasicontinuous if for each $x\in X$, there is a directed family $D(x)\subseteq fin_{n}(x)=\{F:F\ \text{is finite},F\ll_{n}x\}$ such that $D(x)\rightarrow x$.
+A $\mathbf{T_{0}}$ topological space $X$ is called $n$-quasicontinuous if for each $x\in X$, there is a directed family $D(x)\subseteq fin_{n}(x)=\{F:F\ \text{is finite},F\ll_{n}x\}$ such that $D(x)\rightarrow x$.
 Definition 5.43.
-A $T_{0}$ topological space $X$ is called locally hypercompact (or finitary compact), if for any open subsets $U$ and $x\in U$, there exists some $F\subseteq_{f}X$ such that $x\in(\uparrow F)^{\circ}\subseteq\uparrow F\subseteq U$.
-$X$ is called hypercompactly based if for any open subsets $U$ and $x\in U$, there exists some $F\subseteq_{f}X$ such that $x\in(\uparrow F)^{\circ}=\uparrow F\subseteq U$.
-Since $G\ll_{n}H$ iff $H\subseteq(\uparrow G)^{\circ}$, definitions of $n$-quasicontinuous and locally hypercompact are equivalent.
+A $\mathbf{T_{0}}$ topological space $X$ is called locally hypercompact (or finitary compact), if for any open subsets $U$ and $x\in U$, there exists some $F\subseteq_{f}X$ such that $x\in(\uparrow\hskip-2.0ptF)^{\circ}\subseteq\uparrow\hskip-2.0ptF\subseteq U$.
+$X$ is called hypercompactly based if for any open subsets $U$ and $x\in U$, there exists some $F\subseteq_{f}X$ such that $x\in(\uparrow\hskip-2.0ptF)^{\circ}=\uparrow\hskip-2.0ptF\subseteq U$.
+Since $G\ll_{n}H$ iff $H\subseteq(\uparrow\hskip-2.0ptG)^{\circ}$, definitions of $n$-quasicontinuous and locally hypercompact are equivalent.
 Similarly, we can define $n$-quasialgebraic spaces.
@@
 Theorem 5.44.
-[9, 19] For a $T_{0}$ topological space $X$, the following are equivalent:
+[9, 19] For a $\mathbf{T_{0}}$ topological space $X$, the following are equivalent:
   1. (1)
@@
 Definition 5.45.
-A $T_{0}$ topological space is called quasicontinuous if one of the four equivalent conditions in Theorem 5.38 holds; called quasialgebraic if it is $n$-quasialgebraic.
+A $\mathbf{T_{0}}$ topological space is called quasicontinuous if one of the four equivalent conditions in Theorem 5.44 holds; called quasialgebraic if it is $n$-quasialgebraic.
 Next, we show that for a monotone determined space $X$, $\mathcal{O}(X)$ is algebraic iff it is hyperalgebraic.
 Lemma 5.46.
-[4, Lemma 2.3] Let $X$ be a topological space and $U,V\in\mathcal{O}(X)$. $U\prec V$ iff $U\subseteq\uparrow F\subseteq V$ for some $F\subseteq_{f}X$.
+[4, Lemma 2.3] Let $X$ be a topological space and $\in\mathcal{O}(X). UVU\prec V$ iff $U\subseteq\uparrow\hskip-2.0ptF\subseteq V$ for some $F\subseteq_{f}X$.
 Theorem 5.47.
@@
 Remark 5.48.
-Lemma 5.41 does not always hold for general $T_{0}$ spaces.
+Lemma 5.47 does not always hold for general $\mathbf{T_{0}}$ spaces.
 A simple example is $\mathbb{N}$, the set of natural numbers, endowed with the cofinite topology.
 Any open subset of it is compact, but not hypercompact.
-By the above discussion, we conclude the relationships between the continuity of a $T_{0}$ space and the lattice of its open subsets as follows.
+By the above discussion, we conclude the relationships between the continuity of a $\mathbf{T_{0}}$ space and the lattice of its open subsets as follows.
 Theorem 5.49.
-Let $X$ be a $T_{0}$ space.
+Let $X$ be a $\mathbf{T_{0}}$ space.
   1. (1)
@@
     $X$ is quasialgebraic iff $\mathcal{O}(X)$ is a hyperalgebraic lattice. Moreover, if $X$ is a monotone determined space, then $X$ is quasialgebraic iff $\mathcal{O}(X)$ is an algebraic lattice.
-Given any $T_{0}$ space $X$, denote $C(X)$ the lattice of its closed subsets.
-We show that the quasicontinuity of $X$ can be characterized by the quasicontinuity of $C(X)$, and vice versa.
-We first recall some equivalent conditions for continuity of lattices.
+Given any $\mathbf{T_{0}}$ space $X$, denote $C(X)$ the lattice of its closed subsets.
+We will study the relations between $X$ and $C(X)$.
+We first recall some useful conditions for continuity and distributiveness of lattices.
 Lemma 5.50.
@@
     A complete lattice $L$ is a quasicontinuous (quasialgebraic) lattice iff $\omega(L)$ is a continuous (algebraic) lattice.
-Now, it is time to show one of the main results.
 Proposition 5.52.
-Let $X$ be a $T_{0}$ space. If $C(X)$ is a quasicontinuous lattice, then $X$ is quasicontinuous.
+Let $X$ be a $\mathbf{T_{0}}$ space. If $C(X)$ is a continuous lattice, then $X$ is continuous.
+Proof.
+$C(X)$ is a completely distributive lattice by Lemma 5.50 since it has enough co-primes. It follows that $\mathcal{O}(X)$ is completely distributive. Hence $X$ is continuous by Theorem 5.49(1). ∎
+The above result can be generalized to quasicontinuity. Whether an analogous result holds for core-compactness, however, is still unknown.
+Proposition 5.53.
+Let $X$ be a $\mathbf{T_{0}}$ space. If $C(X)$ is a quasicontinuous lattice, then $X$ is quasicontinuous.
 Proof.
 Let $C(X)$ be a quasicontinuous domain.
-Then for any $x\in X$, there exists some $\{A_{1},\dots,A_{n}\}\ll\ \downarrow x$ in $C(X)$.
-Given any $a_{i}\in A_{i}\ (1\leq i\leq n)$, we have $\{\downarrow a_{1},\dots,\downarrow a_{n}\}\ll\ \downarrow x$.
+Then for any $x\in X$, there exists some $\{A_{1},\dots,A_{n}\}\ll\ \downarrow\hskip-2.0ptx$ in $C(X)$.
+Given any $a_{i}\in A_{i}\ (1\leq i\leq n)$, we have $\{\downarrow\hskip-2.0pta_{1},\dots,\downarrow\hskip-2.0pta_{n}\}\ll\ \downarrow\hskip-2.0ptx$.
 Claim 1. $\{a_{1},\dots,a_{n}\}\ll_{n}x$ in $X$.
 Given any net $(x_{i})_{I}\to x$, $x\in\overline{\bigcup_{i\in I}\{x_{i}\}}$.
-Then $\downarrow x\subseteq\ \overline{\bigcup_{i\in I}\{x_{i}\}}$.
-Let =$\mathcal{F}=\{\bigcup_{i\in F}\!\downarrow x_{i}:F\subseteq_{f}I\}$.
-Then $\mathcal{F}$ is directed and $\downarrow x\leq\bigvee^{\uparrow}\mathcal{F}=\overline{\bigcup_{i\in I}\{x_{i}\}}$ in $C(X)$.
-There exist some $F\subseteq_{f}I$ and $a_{j}$ such that $\downarrow a_{j}\subseteq\bigcup_{i\in F}\downarrow x_{i}$.
+Then $\downarrow\hskip-2.0ptx\subseteq\ \overline{\bigcup_{i\in I}\{x_{i}\}}$.
+Let $\mathcal{F}=\{\bigcup_{i\in F}\!\downarrow\hskip-2.0ptx_{i}:F\subseteq_{f}I\}$.
+Then $\mathcal{F}$ is directed and $\downarrow\hskip-2.0ptx\leq\bigvee^{\uparrow}\hskip-2.0pt\mathcal{F}=\overline{\bigcup_{i\in I}\{x_{i}\}}$ in $C(X)$.
+There exist some $F\subseteq_{f}I$ and $a_{j}$ such that $\downarrow\hskip-2.0pta_{j}\subseteq\bigcup_{i\in F}\downarrow\hskip-2.0ptx_{i}$.
 Thus, there exists a $i\in F$ such that $a_{j}\leq x_{i}$, i.e., $\{a_{1},\dots,a_{n}\}\ll_{n}x$ in $X$.
-Claim 2. For any $x$ and any open neighbourhood $U$ of $x$, there exists some $F\subseteq_{f}U$ such that $x\in(\uparrow F)^{\circ}$.
-Given any directed family $\mathcal{A}=\{A_{i}\}_{I}$ of closed subsets of $X$ such that $\downarrow\!x\leq\bigvee\mathcal{A}=\overline{\bigcup_{i\in I}A_{i}}$, then $U\cap(\bigcup_{i\in I}A_{i})\not=\emptyset$.
-There exists some $i\in I$ such that $A_{i}\cap U\not=\emptyset$.
-Pick $b\in A_{i}\cap U$, then $\downarrow\!b\subseteq A_{i}$.
-We have $\{\downarrow\!b:b\in U\}\ll\ \downarrow\!x$ in $C(X)$.
-Since $C(X)$ is quasicontinuous, there exists a $\{B_{1},\dots,B_{m}\}\subseteq C(X)$ such that $\{\downarrow b:b\in U\}\ll\{B_{1},\dots,B_{m}\}\ll\ \downarrow x$ in $C(X)$.
+Claim 2. For any $x$ and any open neighbourhood $U$ of $x$, there exists some $F\subseteq_{f}U$ such that $x\in(\uparrow\hskip-2.0ptF)^{\circ}$.
+Given any directed family $\mathcal{A}=\{A_{i}\}_{I}$ of closed subsets of $X$ such that $\downarrow\hskip-2.0pt\!x\leq\bigvee\mathcal{A}=\overline{\bigcup_{i\in I}A_{i}}$, then $U\cap(\bigcup_{i\in I}A_{i})\not$$\emptyset$.
+There exists some $i\in I$ such that $A_{i}\cap U\not$$\emptyset$.
+Pick $b\in A_{i}\cap U$, then $\downarrow\hskip-2.0pt\!b\subseteq A_{i}$.
+We have $\{\downarrow\hskip-2.0pt\!b:b\in U\}\ll\ \downarrow\hskip-2.0pt\!x$ in $C(X)$.
+Since $C(X)$ is quasicontinuous, there exists a $\{B_{1},\dots,B_{m}\}\subseteq C(X)$ such that $\{\downarrow\hskip-2.0ptb:b\in U\}\ll\{B_{1},\dots,B_{m}\}\ll\ \downarrow\hskip-2.0ptx$ in $C(X)$.
 Then
-$\forall 1\leq i\leq m,\ \{\downarrow b:b\in U\}\ll B_{i}=\overline{\bigcup_{F\subseteq_{f}B_{i}}\downarrow F},$
-which implies that there exists some $F\subseteq_{f}B_{i}$ and $b_{i}\in U$ such that ${\downarrow}b_{i}\subseteq{\downarrow}F\subseteq B_{i}$.
+$\forall 1\leq i\leq m,\ \{\downarrow\hskip-2.0ptb:b\in U\}\ll B_{i}=\overline{\bigcup_{F\subseteq_{f}B_{i}}\downarrow\hskip-2.0ptF},$
+which implies that there exists some $F\subseteq_{f}B_{i}$ and $b_{i}\in U$ such that ${\downarrow\hskip-2.0pt}b_{i}\subseteq{\downarrow\hskip-2.0pt}F\subseteq B_{i}$.
 Thus, $b_{i}\in B_{i}\cap U$.
-It follows that $\{{\downarrow}b_{1},\dots,{\downarrow}b_{m}\}\ll\ {\downarrow}x$ in $C(X)$.
-Then $\{b_{1},\dots,b_{n}\}\ll_{n}x$ by Claim 1, i.e., $x\in({\uparrow}\{b_{1},\dots,b_{n}\})^{\circ}$.
+It follows that $\{{\downarrow\hskip-2.0pt}b_{1},\dots,{\downarrow\hskip-2.0pt}b_{m}\}\ll\ {\downarrow\hskip-2.0pt}x$ in $C(X)$.
+Then $\{b_{1},\dots,b_{n}\}\ll_{n}x$ by Claim 1, i.e., $x\in({\uparrow\hskip-2.0pt}\{b_{1},\dots,b_{n}\})^{\circ}$.
 ∎
 Theorem 5.54.
-Let $X$ be a $T_{0}$ space.
+Let $X$ be a $\mathbf{T_{0}}$ space.
   1. (1)
@@
 ∎
+Finally we collect some equivalent characterizations of quasicontinuous space as follows.
 Corollary 5.55.
-Let $X$ be a $T_{0}$ topological space. Then the following conditions are equivalent:
-  • (1)
+Let $X$ be a $\mathbf{T_{0}}$ topological space. Then the following conditions are equivalent:
+  1. (1)
     $\mathcal{O}(X)$ with the upper topology is corecompact;
-  • (2)
-    $\mathcal{O}(X)$ with the upper topology is an injective space;
-  • (3)
+  2. (2)
+    ([19, Theorem 2(2)]) $\mathcal{O}(X)$ with the upper topology is an injective space;
+  3. (3)
     $C(X)$ is a quasicontinuous domain;
-  • (4)
-    $C(X)$ with Lawson topology is Hausdorff;
-  • (5)
+  4. (4)
+    ([11, Theorem VII-3.10]) $C(X)$ with Lawson topology is Hausdorff;
+  5. (5)
     $X$ is a quasicontinuous space.
-6 Scott completion
-The concept of Scott completion is introduced by Zhang, Shi and Li in [25].
-It has played an important role in constructing free dcpo algebras over general dcpos, see [4, 6].
-But the notion does not seem to have been investigated much.
-We will give more examples of topological spaces which do not have Scott completions.
-Theorem 6.49.
-Let $X$ be a $T_{0}$ space and $(\tilde{X},i)$ be a Scott-completion of $X$, then $\mathcal{O}(X)\cong\mathcal{O}(\tilde{X})$.
-Proof.
-Let $[Y\rightarrow 2]$ denote the set of continuous functions from $Y$ to the Sierpinski space $2$ with the pointwise order.
-It is well known that $\mathcal{O}(Y)\cong[Y\rightarrow 2]$.
-For any $f\in[X\rightarrow 2]$, there exists only one $\bar{f}\in[\tilde{X}\rightarrow 2]$ such that $f=\tilde{f}\circ i$ since $(\tilde{X},i)$ is a Scott-completion of $X$.
-Hence we obtain a map $\theta\colon[X\rightarrow 2]\rightarrow[\tilde{X}\rightarrow 2]$ as follows
-$\theta(f)=\bar{f}.$
-Clearly $\theta$ is bijective.
-$\theta$ is monotone.
-For any $f_{1},f_{2}\in[X\rightarrow 2]$ with $f_{1}\leq f_{2}$.
-We need to show that $\bar{f_{1}}\leq\bar{f_{2}}$.
-Without loss of generality, we consider $y\in\bar{X}$ with $\bar{f_{1}}(y)=1$, i.e., $y\in\bar{f_{1}}^{-1}(1)$.
-We claim that $i(X)\cap{\downarrow}y\cap\bar{f_{1}}^{-1}(1)\neq\emptyset$.
-Otherwise we have the indicator function $\mathbbm{1}_{U}$ of $U=({\downarrow}y)^{c}\cap\bar{f_{1}}^{-1}(1)$ with $\mathbbm{1}_{U}\circ i=\bar{f}_{1}\circ i=f$.
-It follows that $\mathbbm{1}_{U}=\bar{f}_{1}$.
-Then $U=\mathbbm{1}_{U}^{-1}(1)=\bar{f}^{-1}_{1}(1)$.
-This is impossible.
-Back to $i(X)\cap{\downarrow}y\cap\bar{f_{1}}^{-1}(1)\neq\emptyset$.
-We have some $x\in X$ with $i(x)\in{\downarrow}y\cap\bar{f_{1}}^{-1}(1)$.
-It means that $f_{1}(x)=\bar{f}_{1}(i(x))=1$.
-Since $f_{1}\leq f_{2}$, we obtain that $f_{2}(x)=1$.
-Then $\bar{f}_{2}(y)\geq\bar{f}_{2}(i(x))=f_{2}(x)=1$.
-It is easy to see that $\theta$ reflects order. Hence $\theta$ is an order isomorphism. So we obtain that $\mathcal{O}(X)\cong\mathcal{O}(\tilde{X})$. ∎
-Remark 6.50.
-The above proof is similar to the argument of the order isomorphism between the open lattice of a topological space and the open lattice of its soberification.
-The map $i^{-1}\colon\mathcal{O}(\tilde{X})\rightarrow\mathcal{O}(X)$ is an order isomorphic map.
-Furthermore, $i\colon X\rightarrow\tilde{X}$ is an embedding since $i$ is injective, continuous and almost open.
-Corollary 6.51.
-  1. (1)
-    Every $T_{1}$ space endowed with non-discrete topology does not have a Scott completion.
-  2. (2)
-    Every sober but non-Scott space does not have a Scott-completion.
-Proof.
-(1) Assume that there is a $T_{1}$ space $X$ with non-discrete topology such that it has a Scott completion $(\Sigma D,i)$, where $D$ is a dcpo.
-By Theorem 6.49, $\mathcal{O}(X)\cong\sigma(D)$.
-It follows that $C(X)\cong\Gamma(D)$.
-Notice that any singleton set is a minimal element in $C(X)\backslash\{\emptyset\}$.
-Since every minimal nonempty Scott closed of $D$ must be the form of $\{x\}$ ($x\in\text{Min}D$), then $D={\uparrow}\text{Min}D$.
-It is easy to see that $D$ has the discrete order.
-Hence $\Gamma(D)$ is a completely distributive lattice.
-It follows that $X$ is a c-space.
-Hence $X$ is a discrete space.
-Contradiction.
-(2) Assume that there is a sober space $X$ that is not a Scott space and it has a Scott-completion $(\tilde{X},i)$.
-Then $i(X)\subsetneqq\tilde{X}$.
-Consider the soberification map $r\colon\tilde{X}\rightarrow\tilde{X}^{s}$, then $r\circ i\colon X\rightarrow\tilde{X}^{s}$ is also a soberification map.
-This is impossible.
-∎
-The following result is a direct application of Theorem 6.49.
-Corollary 6.52.
-[17, Theorem 4.1] The category of continuous dcpos with Scott continuous maps is not reflective in the category of quasicontinuous dcpos with Scott continuous maps.
-Proof.
-Notice that $(2,\leq)$ is a continuous domain.
-The Sierpinski space is $(2,\leq)$ with its Scott topology.
-Take any quasicontinuous dcpo $D$ which is not continuous.
-Assume that there exists a continuous dcpo $\tilde{D}$ with a Scott continuous map $i\colon D\rightarrow\tilde{D}$, such that it satisfies the universal property.
-By the similar argument in Theorem 6.49, we have $\sigma(D)\cong\sigma(\tilde{D})$.
-This is impossible.
-∎
-Proposition 6.53.
-Let $X$ be a $T_{0}$ space such that it has a Scott-completion $(\tilde{X},i)$. Then $\eta\colon X\rightarrow\Sigma C(X)$ is continuous.
-Proof.
-Notice that $j=\lambda x.{\downarrow}x\colon\tilde{X}\rightarrow\Sigma\Gamma(\tilde{X})$ is continuous since $\tilde{X}$ is a Scott space.
-$i^{-1}\colon\Gamma(\tilde{X})\rightarrow C(X)$ is an order isomorphism by Remark 6.50.
-We claim that
-$\eta=i^{-1}\circ j\circ i.$
-For any $x\in X$, $i^{-1}\circ j\circ i(x)=i^{-1}({\downarrow}i(x))\supseteq{\downarrow}x=\eta(x)$.
-On other hand, for any $y\in i^{-1}({\downarrow}i(x))$, i.e., $i(y)\leq i(x)$.
-Since $i\colon X\rightarrow\tilde{X}$ is an embedding by Remark 6.50, $i$ reflects order.
-It follows that $y\leq x$.
-Therefore $\eta$ is continuous.
-∎
-Corollary 6.54.
-Let $X$ be a $T_{0}$ topological space such that its D-completion is a Scott space. Then $\eta\colon X\rightarrow\Sigma C(X)$ is continuous.
 6 Open questions
@@
   3. (3)
-    Let $X$ be a $T_{0}$ space and $\Sigma C(X)$ be core-compact. Is $X$ core-compact?
+    Let $X$ be a $\mathbf{T_{0}}$ space and $\Sigma C(X)$ be core-compact. Is $X$ core-compact?
 Acknowledgements We thank Yuxu Chen and Peixian Zhou for helpful discussions. We also would like to thank one anonymous referee who points out that one of our main results is not new.